### v1

# Intergenerational Equity of Climate Change Mitigation

Abstract
Today’s decisions on climate change mitigation affect the damage that future generations will bear. Discounting future benefits and costs of climate change mitigation is one of the most important components of assessing efficient climate mitigation pathways. We extend the DICE model with stochastic discount rates to reflect the dynamic nature of discount rates. Stochastic rates give rise to a mitigation strategy, resulting in all model quantities becoming stochastic. We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP and that this effect worsens under stochastic discount rates. Accounting for additional financing risks, we investigate two modifications of DICE. We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing. We propose a modified optimisation to keep costs below 3% of GDP, resulting in more equal efforts between generations.
Introduction
Climate change is one of the most significant risks of the next centuries 1, as demonstrated by its manifold impacts on society 2. Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide3, a transition to (net) zero emissions is necessary to limit global warming to any constant level. A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world. One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see 4; 5.
We focus on the DICE model 6; 7 as a simple IAM, which is used to demonstrate the impacts of modifications to the original model due to its simplicity 8; 9; 10; 11; 12. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points 13; 14; 15; 16; 17; 18; 19, natural feedback processes such as permafrost 9 or abstract catastrophic risks 20; 21. Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimisation function have a strong influence on the resulting optimal pathways 22; 23; 24; 25; 26. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks 13 or inequalities in the distribution of damages 27.
Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation. Model extensions updating IAMs 11; 12; 19; 28 show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality 29 or energy consumption 30, but also general SCC 31 emphasise the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels. Discount rates, in particular, are an essential component of IAM modelling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE 7 versus a normative approach to discounting as in the Stern review 22 continues. The assumptions on discounting have already been discussed in the context of intergenerational 32 and intragenerational inequality 33. Recent studies 34; 35 show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement.
We examine the calibration of the DICE model with respect to intergenerational equity. Measuring the costs per GDP of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretisation 36 and couple this model with a comprehensive modelling toolbox for financial markets 37 to enable the integration of valuation methods for complex financial products in IAMs. We add stochastic interest rates to the model and introduce a stochastic abatement policy. The stochastic abatement policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calibration of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that will significantly improve intergenerational equity: Financing of abatement costs and non-linear discounting of costs, 38. As abatement is a planned investment process, we assume that these costs may be financed by loans with stochastic interest rates. Furthermore, as large projects and investments often incur additional costs 39; 40, we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks. We provide an overview of the extensions in fig. 1. More background is provided in our extensive model description 36.
Basic model DICE-2016-R 7  –  re-implementation with improvements Integrating numerical methods37 (Monte-Carlo simulation, algorithmic differentiation) Improved representation of time-discretisation via an Euler-scheme. Stochastic interest rates g Funding of abatement cost g Non-linear funding of large damage cost Limiting total cost by GDP share
Figure 1: Interest rate related model extensions. We expand our re-implementation of DICE-2016-R with 1) a stochastic interest rate module with an optional stochastic abatement model, 2) the option to allow funding of abatement cost, and, 3) the option to apply non-linear discounting of damage cost. The latter can be used to penalize emission paths that generate cost overruns in terms of cost-per-GDP.
Results
Intergenerational Inequality
The standard DICE model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor.
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (1)
Here $V$ denotes the total welfare, $C_{\mathrm{D}}(t)$ and and $C_{\mathrm{A}}(t)$ denote the damage and abatement cost, respectively, and $T^{\mu=1}$ denotes a parameter representing the abatement policy. The formula presented here is for a reduced model parameterised by the parameter $T^{\mu=1}$, the time by which 100% abatement is achieved.
The criterion eq. 1 is plausible if a single agent optimises its cost but may be problematic if the costs are distributed between different generations. Note that the two individual contributions in eq. 1 occur at different magnitudes at different times, even different generations. fig. 2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The calibration is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
Hence, marginal gains of one generation are equated to marginal losses of another generation. The calibration does not equalise the burden but the weighted marginal burden. In this sense, the calibration likely introduces intergenerational inequality.
Including stochastic interest rates into DICE
We see in eq. 1 that both the welfare cost sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the intertemporal optimal pathway. Including a stochastic rate model, the numeraire ratio becomes stochastic. Allowing for (stochastic) adaptation of the abatement policy to the interest rate level, all model quantities become stochastic.
We show that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality. fig. 3 depicts the impact of stochastic interest rates, both in average and in a quantile measure, on emissions, optimal abatement, damage and total cost-per-GDP. These quantities are depicted for the classical DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10 % and 90 % quantile values.
Figure 3: Classical DICE (grey) and Stochastic DICE Model (green). Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10 % to 90 % percentiles.
Including funding of abatement and funding risks of damage into DICE
We endow the model with two model extensions: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid. Second, we include a non-linear dependency of discounting of the damage cost on the magnitude of the cost. This can be viewed as the additional cost associated with financing risk of very large amount 38.
We find that both modifications allow to improve intergenerational equity (in expectation and risk) in the respective optimal emission pathway. fig. 4 shows the effects of the two modifications on the DICE model with stochastic interest rates. Similar results for the DICE model with deterministic interest rates are shown in LABEL:supfig:pathways_deterministic_with_extension.
Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
Shaded area shows the 10 % to 90 % percentiles.
Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future. fig. 3 shows that the trade-off between early abatement costs and later damages can be balanced by the constraint that total costs are limited to a certain share of GDP. This measure is already used to assess national debt 41 and results in a more equal distribution of damages between generations. Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs. Appropriate parameters of the non-linear discounting of the damage cost can be used to limit the costs per GDP.
To illustrate the unequal burdens between generations, we use population projections from 2015 to 2100 42 to calculate burdens experienced during the projected lifetime. We measure the burden in percentage-of-GDP, where we choose 3 % as the neutral level (green). While differences are small for the time covered by population projections (fig. 5 and LABEL:supfig:generational_burden_deterministic), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3% limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
Figure 5: Illustration of generational inequality: Average cost relative to GDP by year of birth in the different model specifications up to 2200 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
Discussion
We extend the DICE integrated assessment model and compare several extensions for their effect on intergenerational equity. Our analyses apply to integrated assessment models in general and could also be extended to benefit-cost analysis of climate mitigation. We provide a complete open-source implementation of the model illustrating all our extensions, utilising (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other. [^1] The code and data is available at https://gitlab.com/finmath/finmath-climate-nonlinear.
Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP. This is because the calibration does not consider any intergenerational equity but balances the marginal cost changes induced by a policy change. Absolute cost and its distribution over time are irrelevant to the objective function. This effect is driven by the aggregation of discounted welfare into intergenerational welfare. Considering stochastic interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost. This indicates that a repeated re-calibration of the DICE model to current interest rate levels will amplify intergenerational inequality. We use a more applied scenario by considering the funding of abatement cost and non-linear damage costs to show that these additional aspects would increase the incentive for early mitigation, reducing the burden on later generations. More importantly, the non-linear damage cost allow to effectively limit the generational cost to a specific percentage of the GDP.
Using a simple IAM includes several caveats. First, the representation of damages could be improved 43; 11; 12. Second, the lack of regional resolution compared to, e.g., econometric damage estimates 44; 45 ignores regional inequalities in exposure and potential distributional effects of projected costs 46. Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways 31; 47.
Modifying the DICE model in a modular way, we address the issue that the classical model does not consider equity between generations. The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimisation. This is amplified when considering stochastic interest rate risk. Constraining total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
Methods
Derivation of intergenerational inequality in the standard model eq. 1
We consider the one-parameter deterministic abatement model
$\mu(t)\ =\ \min\left(\mu(0)+\frac{1-\mu(0)}{T^{\mu=1}}\ t,1.0\right)\text{,}$
where the parameter $T^{\mu=1}$ represents the time for reaching 100% abatement. We note that the calibration of a model with the full parameter set is close to the calibration of the model reduced to one parameter 36. As the calibration is maximising the total discounted value, we find that the optimal parameter $T^{\mu=1}$, representing the equilibrium state of the model, fulfils
$\frac{\mathrm{d}}{\mathrm{d}T^{\mu=1}}\ \int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t$ $\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t$
[…]
where $C_{\mathrm{A}}(t)$ and $C_{\mathrm{D}}(t)$ denote the abatement cost and damage cost, respectively.
Equation 2 allows us to interpret a change in a parameter of the abatement policy as a change in the cash-flow structure of a financial product. Here $\mathrm{d}t(C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t))$ corresponds to the time-$t$ cash-flow, $\frac{\partial V}{\partial C(t)}$ to a weight (like a notional) and $\frac{N(0)}{N(t)}$ to the discount factor.
From eq. 2, we see that the optimisation across generations may result in intergenerational inequality: Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}T^{\mu=1}>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}T^{\mu=1}<0$, the calibration is equating the marginal cost changes of abatement and damage. However, these changes occur at different times, maybe even at different generations. Note that eq. 2 implies eq. 1.
Details on different extensions of DICE-2016-R
We endowed the classical DICE model with several extensions. These are not specific to the DICE model and can also be applied to other IAMs. They are also independent and parameterised in a way that allows to recover the classical model for a specific parameter choice.
Stochastic Discount Rates
The interest rate $r$ is modelled with a classical interest rate model. The results depicted here were produced with a Hull-White model with a constant mean-reversion and volatility. The implementation allows to replace this model with a general term structure model.
Stochastic Abatement Policy
It is well understood that the optimal abatement policy depends on the interest rate level. Hence, postulating stochastic interest rates, the abatement policy should be allowed to adapt to the interest rate level. This can also be interpreted as illustrating the effect of a repeated re-calibration of the model to changed interest rate levels. We assume a simple, functional model that depends on the spot interest rate, namely
$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot r(t,\omega)\right)\cdot t,1.0\right)\text{.}$ (3)
The simplicity of the models allows us to interpret the parameters. We have that $a_{1}$ in (3) describes the change of the abatement speed by interest rate change,
$a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
Funding of Abatement Cost
Let $C_{\mu}(t)$ denote the cost that occur in time $t$ due to the abatement policy $\mu$. In the classical model, the total cost that occur in time $t$ are
$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
where $C_{\mathrm{D}}(t)$ are the damage costs and $C_{\mathrm{A}}(t)=C_{\mu}(t)$ are the abatement costs.
We alter the model by allowing a funding period, such that abatement costs become effective after a period $\Delta T_{\mathrm{A}}$. For the funding, interest has to be paid such that
$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right).$ (4)
For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, $C_{\mu}=C_{\mathrm{A}}$, and we recover the original model.
Since damages occur instantaneously, we do not consider funding of these. However, we consider a kind of non-linear discounting.
Non-Linear Discounting of Damage Cost
The idea behind non-linear discounting is that the discount factor depends on the level of the amount to be discounted. This can be justified by a model where one has to attribute for the higher default risk of excessively large future cash-flows. The mechanics are that the compensation of this default risk lowers the return (the interest rate) and hence increases the discount factor38.
Put differently, this adds an increasing weight to large or very large damages.
Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then re-define the effective damage costs as
$C_{\mathrm{D}}(t)\ =\ C^{\circ}_{\mathrm{D}}(t)\cdot DC(C^{\circ}_{\mathrm{D}}(t);t)\text{,}$
where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost of funding large projects.
In our experiment, we can specify this weight as a function of the damage $C_{\circ}(t)$ or as a function of the fraction of the GDP, i.e., $C_{\circ}(t)/GDP(t)$.
The latter approach allows us to penalise damages that exceed a certain percentage of the GDP.
Supplementary Material
Supplementary material is available online.
Code and Data availability
The model code for the stochastic DICE model including the proposed extensions, as well as the experiments generating the data and figures is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
Acknowledgements
We thank Max Singhoff for exploring the calibration of the DICE model with the ADAM solver used in our numerical experiments.
L.Q. received funding from the German Federal Ministry of Education and Research (BMBF) under the research projects QUIDIC (01LP1907A) as well as by Deutsche Gesellschaft für Internationale Zusammenarbeit (GIZ) GmbH on behalf of the Government of the Federal Republic of Germany and Federal Ministry for Economic Cooperation and Development (BMZ).
Author Contribution
C.F. and L.Q. developed the research idea together. C.F. lead the re-implementation of the DICE model and extensions as well as the experiments with input from L.Q. L.Q. lead the plotting implementation with input from C.F. C.F. and L.Q. analysed and interpreted results. C.F. derived the analytic expression for generational inequality. C.F and L.Q. wrote the manuscript.
Competing Interests
The authors declare that they have no competing interests.
Supplementary information for
Accounting for Financing Risks improves
Intergenerational Equity of Climate Change Mitigation
Christian P. Fries and Lennart Quante
[…]
figureSupplementary figures
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
supfigure]supfig:pathways_deterministic_with_extension
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
supfigure]supfig:generational_burden_deterministic
Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$. This is the nominal sensitivity to abatement policy changes, entering the calibration.

### v2

@@
-Intergenerational Equity of Climate Change Mitigation
+Positive Effects of Financing
 Abstract
 Today’s decisions on climate change mitigation affect the damage that future generations will bear.
-Discounting future benefits and costs of climate change mitigation is one of the most important components of assessing efficient climate mitigation pathways.
-We extend the DICE model with stochastic discount rates to reflect the dynamic nature of discount rates.
-Stochastic rates give rise to a mitigation strategy, resulting in all model quantities becoming stochastic.
-We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP and that this effect worsens under stochastic discount rates.
-Accounting for additional financing risks, we investigate two modifications of DICE.
+Discounting future benefits and costs of climate change mitigation is one of the most critical components of assessing efficient climate mitigation pathways.
+We extend the DICE model with stochastic discount rates to reflect the uncertain nature of discount rates.
+Stochastic rates give rise to a stochastic mitigation strategy, resulting in all model quantities becoming stochastic.
+We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP.
+Further, we show that considering stochastic discount rates and stochastic abatement policies, which can be interpreted as successive re-calibration, increases intergenerational inequality (and adds additional risks).
+Motivated by this, we consider additional financing risks by investigating two modifications of DICE.
 We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing.
-We propose a modified optimisation to keep costs below 3% of GDP, resulting in more equal efforts between generations.
-Introduction
-Climate change is one of the most significant risks of the next centuries 1, as demonstrated by its manifold impacts on society 2.
-Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide3, a transition to (net) zero emissions is necessary to limit global warming to any constant level.
+To conclude our discussion of options to improve intergenerational equity in an IAM, we propose a modified optimization to keep costs below 3 % of GDP, resulting in more equal distribution of efforts between generations.
+1 Introduction
+Climate change is one of the most significant risks of the next centuries [1], as demonstrated by its manifold impacts on society [2].
+Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide [3], a transition to (net) zero emissions is necessary to limit global warming to any constant level.
 A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world.
-One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see 4; 5.
-We focus on the DICE model 6; 7 as a simple IAM, which is used to demonstrate the impacts of modifications to the original model due to its simplicity 8; 9; 10; 11; 12.
-While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points 13; 14; 15; 16; 17; 18; 19, natural feedback processes such as permafrost 9 or abstract catastrophic risks 20; 21.
-Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimisation function have a strong influence on the resulting optimal pathways 22; 23; 24; 25; 26.
-Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks 13 or inequalities in the distribution of damages 27.
+One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
+We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12].
+While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [13, 14, 15, 16, 17, 18, 19], natural feedback processes such as permafrost [9] or abstract catastrophic risks [20, 21].
+Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimization function have a strong influence on the resulting optimal pathways [22, 23, 24, 25, 26].
+Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [13] or inequalities in the distribution of damages [27].
 Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation.
-Model extensions updating IAMs 11; 12; 19; 28 show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality 29 or energy consumption 30, but also general SCC 31 emphasise the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels.
+Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels.
 Discount rates, in particular, are an essential component of IAM modelling.
-Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE 7 versus a normative approach to discounting as in the Stern review 22 continues.
-The assumptions on discounting have already been discussed in the context of intergenerational 32 and intragenerational inequality 33.
-Recent studies 34; 35 show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement.
+Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues.
+The assumptions on discounting have already been discussed in the context of intergenerational [32] and intragenerational inequality [33].
+Recent studies [34, 35] show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement.
+While our focus lies on the effect of interest rates on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g.
+carbon taxation, [37].
+Parallel to this work, an updated DICE-2023 model was developed and published [38] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments.
+Since the final model remains with a deterministic expected discount rate, our contribution remains independent considering extensions with funding mechanisms and a stochastic discount rate model.
 We examine the calibration of the DICE model with respect to intergenerational equity.
 Measuring the costs per GDP of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations.
-We use an implementation of the DICE model with a flexible time discretisation 36 and couple this model with a comprehensive modelling toolbox for financial markets 37 to enable the integration of valuation methods for complex financial products in IAMs.
+We use an implementation of the DICE model with a flexible time discretization [39] and couple this model with a comprehensive modelling toolbox for financial markets [40] to enable the integration of valuation methods for complex financial products in IAMs.
 We add stochastic interest rates to the model and introduce a stochastic abatement policy.
 The stochastic abatement policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calibration of the model.
-Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that will significantly improve intergenerational equity: Financing of abatement costs and non-linear discounting of costs, 38.
+Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [41].
 As abatement is a planned investment process, we assume that these costs may be financed by loans with stochastic interest rates.
-Furthermore, as large projects and investments often incur additional costs 39; 40, we introduce increasing financing costs for large payments via a non-linear discounting.
+Furthermore, as large projects and investments often incur additional costs [42, 43], we introduce increasing financing costs for large payments via a non-linear discounting.
 These higher costs for high damages can be viewed as (additional) financing risks.
@@
 1.
-More background is provided in our extensive model description 36.
-Basic model DICE-2016-R 7  –  re-implementation with improvements Integrating numerical methods37 (Monte-Carlo simulation, algorithmic differentiation) Improved representation of time-discretisation via an Euler-scheme.
-Stochastic interest rates g Funding of abatement cost g Non-linear funding of large damage cost Limiting total cost by GDP share
+More background is provided in the methods section.
+Basic model DICE-2016-R [7]  –  re-implementation with improvements
+Integrating numerical methods[40] (Monte-Carlo simulation, algorithmic differentiation) Improved representation of time-discretisation via an Euler-scheme.
+Stochastic
+interest rates g
+Funding of
+abatement cost g
+Non-linear funding of
+large damage cost
+Limiting total cost by GDP share
 Figure 1: Interest rate related model extensions.
@@
 The latter can be used to penalize emission paths that generate cost overruns in terms of cost-per-GDP.
-Results
-Intergenerational Inequality
-The standard DICE model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor.
+2 Results
+2.1 Intergenerational Inequality
+To find the optimal mitigation pathway, the standard DICE model equates the increment of the cost of damage $\mathrm{d}C_{\mathrm{D}}(t)$ to the increment of the cost of abatement $\mathrm{d}C_{\mathrm{A}}(t)$, both weighted by the welfare-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor $\frac{N(0)}{N(t)}$,
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (7)
+Considering a simplified 1-parameter abatement model parameterizing the abatement policy by $T^{\mu=1}$, the time by which 100% abatement is achieved, this simplifies to
 $\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (9)
-Here $V$ denotes the total welfare, $C_{\mathrm{D}}(t)$ and and $C_{\mathrm{A}}(t)$ denote the damage and abatement cost, respectively, and $T^{\mu=1}$ denotes a parameter representing the abatement policy.
-The formula presented here is for a reduced model parameterised by the parameter $T^{\mu=1}$, the time by which 100% abatement is achieved.
-The criterion eq.
-1 is plausible if a single agent optimises its cost but may be problematic if the costs are distributed between different generations.
+The conditions (7) and (9) are plausible if a single agent optimizes its cost but may be problematic if the costs are distributed between different generations.
 Note that the two individual contributions in eq.
@@
 Hence, marginal gains of one generation are equated to marginal losses of another generation.
-The calibration does not equalise the burden but the weighted marginal burden.
-In this sense, the calibration likely introduces intergenerational inequality.
-Including stochastic interest rates into DICE
-We see in eq.
-1 that both the welfare cost sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the intertemporal optimal pathway.
-Including a stochastic rate model, the numeraire ratio becomes stochastic.
-Allowing for (stochastic) adaptation of the abatement policy to the interest rate level, all model quantities become stochastic.
-We show that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality.
+The calibration does not equalize the burden but the weighted marginal burden.
+In this sense, the calibration is indifferent to the temporal distribution of the absolute burden.
+This may generate intergenerational inequality.
+2.2 Including stochastic interest rates into DICE
+We see in Equation 7 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the inter-temporal optimal pathway.
+Including a stochastic interest rate model, the numéraire $N$ becomes stochastic.
+Allowing for (stochastic) adaptation of the abatement policy $\mu$ to the interest rate level, all model quantities become stochastic.
+We use a classical Hull-White model - details are provided in the methods (11).
+The model generates the same expected discount rate as for the deterministic DICE model we compare it with, which implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model.
+As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [39].
+However, allowing for a stochastic abatement policy that adapts to the interest rate level, e.g., under the model (12) presented in the methods, we see an adverse effect: the expected abatement is smaller, the expected emissions are higher and the expected total cost increase significantly for later years.
+Note that restricting the stochastic abatement model (12) to the case $a_{1}=0$ agrees with the abatement model (4).
+The effect is hence solely created by allowing the change of abatement speed by interest rate change $a_{1}$ to be a free parameter.
 fig.
-3 depicts the impact of stochastic interest rates, both in average and in a quantile measure, on emissions, optimal abatement, damage and total cost-per-GDP.
-These quantities are depicted for the classical DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10 % and 90 % quantile values.
-Figure 3: Classical DICE (grey) and Stochastic DICE Model (green).
+3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure, on emissions, optimal abatement, damage and total cost-per-GDP.
+These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
+Figure 3: Classical DICE (grey), Classical DICE with reduce Abatement Model (4) (blue) and Stochastic DICE Model (green).
+Restricting the abatement policy to the 1-parameter model (4) has a comparably small effect on the emission pathways.
 Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost).
-Shaded area shows the 10 % to 90 % percentiles.
-Including funding of abatement and funding risks of damage into DICE
-We endow the model with two model extensions: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid.
+Shaded area shows the 10th to 90th percentiles.
+This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles).
+It also lowers the abatement and increases the emission.
+Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (10) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates.
+Note that for a convex function $f$ and a random variable $X$ we have $\mathrm{E}(f(X))\leq f(\mathrm{E}(X))$ (Jensen’s inequality).
+2.3 Including funding of abatement and funding risks of damage into DICE
+As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future.
+We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid.
 Second, we include a non-linear dependency of discounting of the damage cost on the magnitude of the cost.
-This can be viewed as the additional cost associated with financing risk of very large amount 38.
-We find that both modifications allow to improve intergenerational equity (in expectation and risk) in the respective optimal emission pathway.
-fig.
-4 shows the effects of the two modifications on the DICE model with stochastic interest rates.
-Similar results for the DICE model with deterministic interest rates are shown in LABEL:supfig:pathways_deterministic_with_extension.
-Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
-Shaded area shows the 10 % to 90 % percentiles.
-Figure 4: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
-As shown in fig.
-3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future.
-fig.
-3 shows that the trade-off between early abatement costs and later damages can be balanced by the constraint that total costs are limited to a certain share of GDP.
-This measure is already used to assess national debt 41 and results in a more equal distribution of damages between generations.
+This can be viewed as the additional cost associated with financing risk of very large amounts [41].
+Details on this approach and parameters are provided in the methods section 4.3.4.
+We find that both modifications improve intergenerational equity in the classical deterministic DICE model, fig.
+4, and in the DICE model with stochastic interest rate, fig.
+5.
+In the latter, intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
+Figure 4: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
+Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
+Shaded area shows the 10th to 90th percentiles.
 Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs.
-Appropriate parameters of the non-linear discounting of the damage cost can be used to limit the costs per GDP.
-To illustrate the unequal burdens between generations, we use population projections from 2015 to 2100 42 to calculate burdens experienced during the projected lifetime.
-We measure the burden in percentage-of-GDP, where we choose 3 % as the neutral level (green).
+Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$).
+For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [48] to calculate burdens experienced during the projected lifetime.
+We measure the burden in percentage-of-GDP (section 4.2.3), where we choose 3 % as the neutral level (green).
 While differences are small for the time covered by population projections (fig.
-5 and LABEL:supfig:generational_burden_deterministic), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century.
+6 and fig.
+7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century.
 Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset.
 The deterministic model keeps the 3% limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
-Figure 5: Illustration of generational inequality: Average cost relative to GDP by year of birth in the different model specifications up to 2200 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
-Discussion
+Figure 6: Illustration of generational inequality, deterministic model: Lifetime-average cost relative to GDP by year of birth in the different model specifications up to 2200.
+Normalized by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
+Figure 7: Illustration of generational inequality, stochastic model: Lifetime-average cost relative to GDP by year of birth in the different model specifications up to 2200.
+Normalized by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
+3 Discussion
+We analyze the DICE integrated assessment model and derive metrics to assess intergenerational equity of the optimal emission pathways calibrated by the model.
 We extend the DICE integrated assessment model and compare several extensions for their effect on intergenerational equity.
 Our analyses apply to integrated assessment models in general and could also be extended to benefit-cost analysis of climate mitigation.
-We provide a complete open-source implementation of the model illustrating all our extensions, utilising (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other.
-[^1] The code and data is available at https://gitlab.com/finmath/finmath-climate-nonlinear.
+We provide a complete open-source implementation of the model illustrating all our extensions, utilizing Monte-Carlo implementation of standard interest rate models for simulation and (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other.
 Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP.
-This is because the calibration does not consider any intergenerational equity but balances the marginal cost changes induced by a policy change.
+This is because the calibration does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change.
 Absolute cost and its distribution over time are irrelevant to the objective function.
@@
 This indicates that a repeated re-calibration of the DICE model to current interest rate levels will amplify intergenerational inequality.
-We use a more applied scenario by considering the funding of abatement cost and non-linear damage costs to show that these additional aspects would increase the incentive for early mitigation, reducing the burden on later generations.
-More importantly, the non-linear damage cost allow to effectively limit the generational cost to a specific percentage of the GDP.
+We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost.
+Funding of abatement cost may be considered as modelling the backing of abatement costs by government loans.
+We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost.
+The non-linear discounting of damage cost may be considered as modelling financing risk of damage costs, where very large cost have higher financing cost.
+The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
 Using a simple IAM includes several caveats.
-First, the representation of damages could be improved 43; 11; 12.
-Second, the lack of regional resolution compared to, e.g., econometric damage estimates 44; 45 ignores regional inequalities in exposure and potential distributional effects of projected costs 46.
-Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways 31; 47.
+First, the representation of damages could be improved [49, 11, 12].
+Second, the lack of regional resolution compared to, e.g., econometric damage estimates [50, 51] ignores regional inequalities in exposure and potential distributional effects of projected costs [52].
+Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways [31, 53].
 Modifying the DICE model in a modular way, we address the issue that the classical model does not consider equity between generations.
-The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimisation.
+The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimization.
 This is amplified when considering stochastic interest rate risk.
-Constraining total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
-Methods
-Derivation of intergenerational inequality in the standard model eq. 1
-We consider the one-parameter deterministic abatement model
-$\mu(t)\ =\ \min\left(\mu(0)+\frac{1-\mu(0)}{T^{\mu=1}}\ t,1.0\right)\text{,}$
+Constraining the total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
+4 Methods
+The starting point of our investigations and extensions is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
+Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model.
+For the equi-distant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered.
+Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$.
+The differences to the classical DICE-2016 are negligible for our work.
+Except where otherwise stated, we use the same model components (e.g.
+abatement cost and damage cost model) as in DICE-2016.
+A detailed description of our implementation of the DICE model (and its extensions) can be found in [39].
+We use the same notation, except that we denote the cost functions with $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
+We will recall the calibration procedure first in Section 4.1, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
+We then present metrics to analyse the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.2, since the question at which time cost occur is a question of inter-generational equity.
+We conclude presenting different extensions to the model, which change the distribution of cost.
+We discuss these in Section 4.3.
+We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
+4.1 Model Calibration
+The classical objective function of the DICE model is the integrated discounted welfare
+$\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\text{.}$ (1)
+Here $N$ denotes the numéraire.
+For the classical model, this means that $N(0)/N(t)$ is the discount factor.
+We use the more general notion of a numéraire, as we will consider a more general model later.
+DICE-2016 is recovered by $N(t)=\exp(r\cdot t)$.
+4.1.1 Objective Function for the Stochastic Model
+In the following we will consider an extension of the model where $N$ and $V$ are stochastic processes. This requires that we re-define the objective function as (1) is a random variable.
+For the calibration we consider the expectation operator
+$\mathrm{E}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{,}$ (2)
+i.e., the objective function for the stochastic model is the expected welfare.
+For analysis purposes we will plot a risk measure, here the value-at-risk risk, i.e., a percentile of the stochastic welfare,
+$\mathrm{VaR}_{\alpha}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{.}$ (3)
+4.1.2 Abatement Policy - Classical DICE Model
+The model calibration determines the optimal abatement policy $t\mapsto\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (1), (2).
+In the time-discrete classical DICE model the calibration determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t)\right)$.
+For a model time horizon of $T=500\leavevmode\nobreak\ \text{years}$ and a time-discretization step of $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$ this results in 499 free parameters for the abatement policy (and another corresponding set for the savings rate).
+4.1.3 Parametric Abatement Policy - Reduced Model
+For analysis purposes and for comparison with a parametric stochastic abatement model, which we will introduce as part of our model extensions later, we define a parametric abatement model
+$\mu(t)\ =\ \min\left(\mu(0)+\frac{1-\mu(0)}{T^{\mu=1}}\ t,1.0\right)\text{,}$ (4)
 where the parameter $T^{\mu=1}$ represents the time for reaching 100% abatement.
-We note that the calibration of a model with the full parameter set is close to the calibration of the model reduced to one parameter 36.
-As the calibration is maximising the total discounted value, we find that the optimal parameter $T^{\mu=1}$, representing the equilibrium state of the model, fulfils
-$\frac{\mathrm{d}}{\mathrm{d}T^{\mu=1}}\ \int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t$ $\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t$
+This model has only a single parameter, however it turns out that the optimal abatement policy in the full parameter model is comparably close to a functional form of the type (4).
+4.2 Cost
+To characterize the resulting optimal mitigation paths, we utilise derived quantities.
+A prominent example of such a derived quantity is the social cost of carbon, $SCC(t)$.
+The $SCC$ is the time-$t$ marginal cost of emitting one additional unit of carbon.
+Its unit is $[SCC]=\frac{\mathrm{USD}}{\mathrm{tCO}_{2}}$.
+The “social cost of carbon” is maybe not a suitable measure to access the temporal distribution of the social burden, and hence intergenerational equity, because pricing emissions at the level of the SCC does not cover the cost associate with the climate mitigation paths: the accumulated value $\int_{0}^{T}SCC(t)\cdot E(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$ does not match the accumulated cost $\int_{0}^{T}C(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$, see [44].
+4.2.1 Distribution of Cost
+We analyse the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
+$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
+To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the calibration’s objective function.
+4.2.2 Sensitivity of Damage and Abatement Cost to Policy Changes
+Differentiating the objective function (1) with respect to the abatement policy $\mu$ gives
+$\frac{\mathrm{d}}{\mathrm{d}\mu}\ \int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t$ $\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C(t)}{\mathrm{d}\mu}\ \frac{N(0)}{N(t)}\mathrm{d}t$
+$\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\left(\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu}+\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu}\right)\ \frac{N(0)}{N(t)}\mathrm{d}t\text{.}$ (5)
+As the objective function depends on the cost, two transformations take place: First, the value $V$ is defined as utility.
+Since the utility function is convex, there is some saturation, resulting in a over time decaying weight $\frac{\partial V(t)}{\partial C(t)}$ applied to the cost.
+Second, the value is discounted.
+For the model of a constant positive discount rate, the discount factor constitutes an exponentially decaying weight $\frac{N(0)}{N(t)}$ applied to the cost.
+The product of both weights represents the cost-to-value-weight.
+The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfils
 […]
-where $C_{\mathrm{A}}(t)$ and $C_{\mathrm{D}}(t)$ denote the abatement cost and damage cost, respectively.
-Equation 2 allows us to interpret a change in a parameter of the abatement policy as a change in the cash-flow structure of a financial product.
-Here $\mathrm{d}t(C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t))$ corresponds to the time-$t$ cash-flow, $\frac{\partial V}{\partial C(t)}$ to a weight (like a notional) and $\frac{N(0)}{N(t)}$ to the discount factor.
 From eq.
-2, we see that the optimisation across generations may result in intergenerational inequality: Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}T^{\mu=1}>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}T^{\mu=1}<0$, the calibration is equating the marginal cost changes of abatement and damage.
-However, these changes occur at different times, maybe even at different generations.
-Note that eq.
-2 implies eq.
-1.
-Details on different extensions of DICE-2016-R
-We endowed the classical DICE model with several extensions.
-These are not specific to the DICE model and can also be applied to other IAMs.
-They are also independent and parameterised in a way that allows to recover the classical model for a specific parameter choice.
-Stochastic Discount Rates
-The interest rate $r$ is modelled with a classical interest rate model.
-The results depicted here were produced with a Hull-White model with a constant mean-reversion and volatility.
-The implementation allows to replace this model with a general term structure model.
-Stochastic Abatement Policy
-It is well understood that the optimal abatement policy depends on the interest rate level.
-Hence, postulating stochastic interest rates, the abatement policy should be allowed to adapt to the interest rate level.
-This can also be interpreted as illustrating the effect of a repeated re-calibration of the model to changed interest rate levels.
-We assume a simple, functional model that depends on the spot interest rate, namely
+6 we see why the model calibration may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered.
+Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq.
+6 that the model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$.
+We have
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{,}$ (7)
+where the to cost function $t\mapsto\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}$ and $t\mapsto\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}$ have different temporal distributions.
+A change in the abatement policy in time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for possibly all $t\geq s$.
+This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$.
+As the model calibration is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity.
+The minus in (8) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
+$\rho_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (8)
+From Equation 5 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density.
+Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
+Figure 8: The density $t\mapsto\rho_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
+In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement (Figure 8), indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
+Sensitivity of Cost to Time of Reaching 100 % Abatement
+The sensitivity $\frac{\mathrm{d}}{\mathrm{d}\mu}$ is a functional derivative or gradient.
+To visualize the aggregated dependency structure, we consider the one-parameter deterministic abatement model (4).
+For this abatement model, the equilibrium state is given by
+[…]
+Given the one-parameter model, we may depict the (weighted) damage cost sensitivity $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ and the (weighted) abatement cost sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$, see fig.
+2.
+4.2.3 Cost per GDP
+To measure the burden born by a generation, we consider the relative total damage and abatement cost per GDP.
+Putting nominal values relative to the GDP is a common measure, e.g., to assess national debt [45].
+Here, it reflects that a more wealthy generation can carry a greater burden, mimicking an effort sharing scheme and also considering the greater capacity of future generations under the expected baseline growth in DICE.
+In our numerical experiments we measure the instantaneous cost-per-GDP $C(s)/GDP(s)$ as well as the lifetime-average cost-per-GDP $\overline{C}(s)/\overline{GDP}(s)$.
+To allow for a netting inside a generation one may consider the (discounted) average value over a generation lifetime.
+For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation.
+We then consider the lifetime average cost and GDP.
+$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
+4.3 Model Extensions
+We present modifications to integrated assessment models to investigate the effect of stochastic interest rates, funding of abatement costs, and non-linear financing cost [41].
+These may be also incorporated into more complex IAMs.
+We use the simple DICE model to illustrate potential effects.
+4.3.1 Stochastic Interest Rates
+The classical way in which interest rates are modelled in an IAM is via a discount factor.
+A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
+$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
+We model stochastic interest rates $r$.
+Our implementation allows to use a general discrete forward rate model (LIBOR Market Model).
+Our results were conducted with a classical Hull-White model.
+The model provides the stochastic numeráire $N$ (and thus discount factor) as well as the stochastic forward rate $FR$.
+Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates).
+For exemplary parameterization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$.
+While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively.
+For a corresponding visualization, see [39].
+If interest rates are stochastic but the function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM since
+$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (10)
+with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\right)$.
+Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics.
+This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process.
+[^2] Our choice to make interest rates stochastic is only exemplary.
+The interest rate is an important factor in linking present abatement cost to the avoided future damage cost.
+While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
+We use a classical Hull-White model, that is
+$\begin{split}\mathrm{d}r(t)&\ =\ (\theta(t)-a(t)r(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad r(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ r(t)\ \mathrm{d}t\text{,}\end{split}$ (11)
+where an exact time-discretization scheme [47] was used.
+The model’s mean-reversion level $\theta$ is such that the numéraire-relative zero coupon bond prices are martingales and their expectation agrees with the classical non-stochastic model.
+In other words, our stochastic interest rate model is such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
+4.3.2 Stochastic Abatement Policy
+Assuming that a change in the interest rates triggers a re-calibration of the model, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic.
+The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates.
+This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
+The determination of an optimal stochastic abatement policy is a classical optimal exercise problem, as in the valuation of financial products with early exercise options in mathematical finance.
+Here, $\mu$ takes the role of the exercise strategy.
+As the time-$t$ optimal exercise is required to be a $\mathcal{F}_{t}$-measurable, it may be represented as a functional form of time-$t$ measurable random variables, see [46].
+An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization.
+As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $r(t)$.
+A simple example of a stochastic abatement model is a parametric one, where the abatement speed is a (linear) function of the interest rate level $r$, e.g.,
 $\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot r(t,\omega)\right)\cdot t,1.0\right)\text{.}$ (12)
-The simplicity of the models allows us to interpret the parameters. We have that $a_{1}$ in (3) describes the change of the abatement speed by interest rate change,
-$a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
-Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
-Funding of Abatement Cost
-Let $C_{\mu}(t)$ denote the cost that occur in time $t$ due to the abatement policy $\mu$. In the classical model, the total cost that occur in time $t$ are
-$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
-where $C_{\mathrm{D}}(t)$ are the damage costs and $C_{\mathrm{A}}(t)=C_{\mu}(t)$ are the abatement costs.
-We alter the model by allowing a funding period, such that abatement costs become effective after a period $\Delta T_{\mathrm{A}}$. For the funding, interest has to be paid such that
-$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right).$ (4)
-For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, $C_{\mu}=C_{\mathrm{A}}$, and we recover the original model.
-Since damages occur instantaneously, we do not consider funding of these. However, we consider a kind of non-linear discounting.
-Non-Linear Discounting of Damage Cost
-The idea behind non-linear discounting is that the discount factor depends on the level of the amount to be discounted.
-This can be justified by a model where one has to attribute for the higher default risk of excessively large future cash-flows.
-The mechanics are that the compensation of this default risk lowers the return (the interest rate) and hence increases the discount factor38.
-Put differently, this adds an increasing weight to large or very large damages.
-Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e. the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$. We then re-define the effective damage costs as
+The simplicity of the models allows us to interpret the parameters.
+We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
+$a_{0}\ =\ \frac{\partial\mu(t)}{\partial t}\big{|}_{r=0}\text{,}\qquad a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
+Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (4) simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
+The calibration of the optimal model (12) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
+4.3.3 Funding of Abatement Costs
+The geophysical part of the model models a global mean temperature level, which results from past emissions.
+The simple damage function of the DICE model provides damages as a reduction of global GDP.
+Damages may be reduced by performing abatement of emissions.
+Both abatement and damage are associated with costs.
+The time $t_{i}$ costs are deduced from the time $t_{i}$ GDP.
+The remainder is then available for consumption or investment.
+Investment adds to the capital which determines the GDP of the next time $t_{i+1}$.
+However, abatement cost and damage cost linked to the same abatement policy decision occur at significantly different times.
+We introduce the option of funding of abatement cost.
+As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest rate corresponds to the current discount rate.
+Our model allows to apply a funding spread, i.
+e.
+higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon.
+In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, as introduced in DICE-2023 [38].
+We define $C_{\mu}(t)$ as the instantaneous abatement costs in time $t$, i.e., the quantity that was formerly denoted by $C_{\mathrm{A}}(t)$.
+We allow that $C_{\mu}(t)$ might be funded for a period $\Delta T_{\mathrm{A}}$.
+Thus abatement costs of $C_{\mu}(t)$ are accrued with the forward rate $FR(t,t+\Delta T_{\mathrm{A}};t)$ observed in $t$, such that the realized abatement cost $C_{\mathrm{A}}$ at time $t+\Delta T_{\mathrm{A}}$ are
+$C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right)\text{.}$ (13)
+For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, we recover the classical model with $C_{\mu}=C_{\mathrm{A}}$.
+In a standard model for risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting.
+However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated and due to the time-preference included in the utility function.
+Since damages occur instantaneously, we do not consider funding of these. The total cost is given by $C(t):=C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)$.
+4.3.4 Non-Linear Financing Costs
+For an unsecured financial cash-flow its present value is defined by a discount factor times the cash-flow.
+If the future cash-flow is subject to default, the discount factor is lower, reflecting the additional value reduction due to the risk of (partial) default.
+As default is not an option for future damage cost, it appears as if the risk-free discount factor should apply.
+However, since no hedging strategy exists, a risk free funding is not possible.
+Thus additional cost may occur to secure the unsecured funding [41].
+Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger cost.
+We model this by optionally adding non-linear financing cost using a non-linear funding model [41].
+This model allows, that the discount factor may depend on the magnitude of the cash-flow.
+Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions.
+For our application, this means that larger cost get a larger weight.
+Our model modification is now a modification of the damage cost.
+Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e.
+the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$.
+We then re-define the effective damage costs as
 $C_{\mathrm{D}}(t)\ =\ C^{\circ}_{\mathrm{D}}(t)\cdot DC(C^{\circ}_{\mathrm{D}}(t);t)\text{,}$
-where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost of funding large projects.
-In our experiment, we can specify this weight as a function of the damage $C_{\circ}(t)$ or as a function of the fraction of the GDP, i.e., $C_{\circ}(t)/GDP(t)$.
-The latter approach allows us to penalise damages that exceed a certain percentage of the GDP.
+where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost to fund large projects.
+As $C^{\circ}_{\mathrm{D}}(t)$ represents a time-value, it is natural that the default compensation factor depends on a normalized value only.
+We allow two different normalizations: either with the numéraire $N(t)$ or with the GDP $Y(t)$.
+Hence, our model for the default compensator is
+$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{N}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{N(t)})\text{,}$ (14)
+or, alternatively,
+$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{Y}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{Y(t)})\text{.}$ (15)
+The latter approach allows us to penalize damages that exceed a certain percentage of the GDP.
+The factor $DC^{*}(x)$ depends on the size of $x$. For small $x$ we have $DC^{*}(x)=1$, but for large $x$ we may have factors $>1$. Obviously, this will penalize large spikes in the costs.
+In our numerical experiments the funding period was set to 20 years and the non-linear discounting was set to increase the default compensator significantly for damage cost over 3 % of the GDP (black).
+The parameter choice is exemplary, other experiments with similar results can be reproduced in the published source code.
+As the abatement cost are usually smaller and part of a more planned process, we do not consider a default compensation factor for the abatement cost.
 Supplementary Material
@@
 Code and Data availability
-The model code for the stochastic DICE model including the proposed extensions, as well as the experiments generating the data and figures is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
+The model code for the stochastic DICE model, including the proposed extensions, as well as the experiments generating the data and figures, is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
 Acknowledgements
@@
 C.F.
-lead the re-implementation of the DICE model and extensions as well as the experiments with input from L.Q.
+lead the re-implementation of the DICE model and extensions, as well as the experiments with input from L.Q.
 L.Q.
@@
 and L.Q.
-analysed and interpreted results.
+analyzed and interpreted results.
 C.F.
@@
 Supplementary information for
-Accounting for Financing Risks improves
-Intergenerational Equity of Climate Change Mitigation
-Christian P. Fries and Lennart Quante
+Intergenerational Equitable Climate Change Mitigation:
+Negative Effects of Stochastic Interest Rates;
+Positive Effects of Financing
+[…]
 […]
@@
 figureSupplementary figures
-Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$.
-This is the nominal sensitivity to abatement policy changes, entering the calibration.
-Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
-supfigure]supfig:pathways_deterministic_with_extension
-Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
-supfigure]supfig:generational_burden_deterministic
-Supplementary Figure 3: Deterministic model: Total cost relative to GDP by year of birth in the different model specifications up to 2100 Normalised by annual GDP and scaled such that blue means lower, red higher average total cost values than 3% of GDP for this generation.
-Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost
-Supplementary Figure 1: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$, without weighting the values by $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{N(0)}{N(t)}$.
-This is the nominal sensitivity to abatement policy changes, entering the calibration.
+Supplementary Figure 1: Increasing interest rate level increases the time of 100% abatement $T^{\mu=1}$.
+Supplementary Figure 2: Increasing interest rate volatility, i.e. riskiness, decreases the time of 100% abatement $T^{\mu=1}$.

### v3

@@
 Stochastic rates give rise to a stochastic mitigation strategy, resulting in all model quantities becoming stochastic.
-We show that the classical calibration of the DICE model induces intergenerational inequality: future generations have to bear higher costs relative to GDP.
-Further, we show that considering stochastic discount rates and stochastic abatement policies, which can be interpreted as successive re-calibration, increases intergenerational inequality (and adds additional risks).
+We show that the optimization procedure of the DICE model induces intergenerational inequality: lacking a mechanism to regulate burden, future generations have to bear higher costs from abatement and damage relative to GDP.
+Further, we show that considering uncertainty of discount rates and their feedback to abatement policies, which can be interpreted as successive re-calculation, increases intergenerational inequality (and adds additional risks).
 Motivated by this, we consider additional financing risks by investigating two modifications of DICE.
@@
 A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world.
-One traditional method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
+One well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
 We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12].
@@
 Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels.
-Discount rates, in particular, are an essential component of IAM modelling.
+Discount rates, in particular, are an essential component of IAM modeling.
 Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues.
 The assumptions on discounting have already been discussed in the context of intergenerational [32] and intragenerational inequality [33].
-Recent studies [34, 35] show that dynamic modelling of discount rates might allow a more precise assessment of the costs of damages and abatement.
-While our focus lies on the effect of interest rates on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g.
+Recent studies [34, 35] show that dynamic modeling of discount rates might allow a more precise assessment of the costs of damage and abatement.
+While our focus lies on the effect of uncertainty about discount rates and stochastic adaptation of abatement policies on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g.
 carbon taxation, [37].
 Parallel to this work, an updated DICE-2023 model was developed and published [38] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments.
-Since the final model remains with a deterministic expected discount rate, our contribution remains independent considering extensions with funding mechanisms and a stochastic discount rate model.
-We examine the calibration of the DICE model with respect to intergenerational equity.
-Measuring the costs per GDP of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations.
-We use an implementation of the DICE model with a flexible time discretization [39] and couple this model with a comprehensive modelling toolbox for financial markets [40] to enable the integration of valuation methods for complex financial products in IAMs.
-We add stochastic interest rates to the model and introduce a stochastic abatement policy.
-The stochastic abatement policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calibration of the model.
+These uncertainties are translated into certainty-equivalent discount rate, such that the overall model remains deterministic.
+As we show in eq.
+11, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
+Complementary to the potential uncertainties in future consumption growth and riskiness of climate investments, we consider uncertainty about the rate of time preference $\rho$ itself.
+For this, we use a standard interest-rate model - the Hull-White model [39] to model stochastic fluctuation of the rate of time preference $\rho$, interacting with an abatement policy adjusting to this uncertainty.
+1.1 Extensions to the DICE Model
+We examine the optimization of the DICE model with respect to uncertainty in the rate of time-preference, and its consequences for intergenerational equity.
+Measuring the costs per GDP (section 4.4.5) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations.
+We use an implementation of the DICE model with a flexible time discretization [40] and couple this model with a comprehensive modeling toolbox for financial markets [41] to enable the integration of valuation methods for complex financial products in IAMs.
+1.2 Interest Rates
+We add stochastic interest rates - through modeling the rate of time preference $\rho$ as stochastic - to the model, and introduce a stochastic abatement policy.
+This policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calculation of the model.
 Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [42].
-As abatement is a planned investment process, we assume that these costs may be financed by loans with stochastic interest rates.
+As abatement is a planned investment process, we assume that these costs may be financed by loans with uncertain interest rates.
 Furthermore, as large projects and investments often incur additional costs [43, 44], we introduce increasing financing costs for large payments via a non-linear discounting.
@@
 Note that the two individual contributions in eq.
-9 occur at different magnitudes at different times, even different generations.
+10 occur at different times.
 fig.
@@
 Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$.
-The calibration is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
+The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
 Hence, marginal gains of one generation are equated to marginal losses of another generation.
-The calibration does not equalize the burden but the weighted marginal burden.
-In this sense, the calibration is indifferent to the temporal distribution of the absolute burden.
+The optimization does not equalize the burden but the weighted marginal burden.
+In this sense, the optimization is indifferent to the temporal distribution of the absolute burden.
 This may generate intergenerational inequality.
 2.2 Including stochastic interest rates into DICE
-We see in Equation 7 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the numeraire ratio $\frac{N(0)}{N(t)}$ (i.e., the discount factor) are important to determine the inter-temporal optimal pathway.
+We see in Equation 8 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the discount factor $\frac{N(0)}{N(t)}$ (i.e., the numeraire ratio) are important to determine the inter-temporal optimal pathway.
 Including a stochastic interest rate model, the numéraire $N$ becomes stochastic.
 Allowing for (stochastic) adaptation of the abatement policy $\mu$ to the interest rate level, all model quantities become stochastic.
-We use a classical Hull-White model - details are provided in the methods (11).
-The model generates the same expected discount rate as for the deterministic DICE model we compare it with, which implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model.
+We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$ details are provided in the methods (12).
+The model generates the same expected discount rate as for the deterministic DICE model.
+This implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model.
 As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [40].
@@
 fig.
-3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure, on emissions, optimal abatement, damage and total cost-per-GDP.
+3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage and total cost-per-GDP.
 These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
@@
 This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles).
-It also lowers the abatement and increases the emission.
+This is qualitatively independent of the interest rate level.
 Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (11) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates.
-Note that for a convex function $f$ and a random variable $X$ we have $\mathrm{E}(f(X))\leq f(\mathrm{E}(X))$ (Jensen’s inequality).
+Note that for a convex function $f$ and a random variable $X$ Jensen’s inequality implies $f(\mathrm{E}(X))\leq\mathrm{E}(f(X))$.
+The strength of the effect increases with increasing interest rate volatility.
+A corresponding analysis for different volatility levels can be found in Variation of the Interest Rate Volatility of the supplementary information.
+For the limits cast of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
 2.3 Including funding of abatement and funding risks of damage into DICE
 As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future.
-We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time for interest that has to be paid.
-Second, we include a non-linear dependency of discounting of the damage cost on the magnitude of the cost.
+We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time, resulting in interest that has to be paid.
+Second, we include a non-linear dependency of the discount factor for the damage cost on the magnitude of the cost.
 This can be viewed as the additional cost associated with financing risk of very large amounts [42].
-Details on this approach and parameters are provided in the methods section 4.3.4.
-We find that both modifications improve intergenerational equity in the classical deterministic DICE model, fig.
+Details on this extension and parameters are provided in the methods section 4.4.4.
+We find that both extensions improve intergenerational equity in the classical deterministic DICE model, fig.
 4, and in the DICE model with stochastic interest rate, fig.
@@
 Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
-Shaded area shows the 10th to 90th percentiles.
+Shaded area shows the 10th to 90th percentiles (VaR).
 Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs.
 Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$).
-For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [48] to calculate burdens experienced during the projected lifetime.
-We measure the burden in percentage-of-GDP (section 4.2.3), where we choose 3 % as the neutral level (green).
+For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [45] to calculate burdens experienced during each birth-year cohorts projected lifetime.
+We measure the burden in percentage-of-GDP (section 4.4.5), where we choose 3% as the neutral level (green), [46].
 While differences are small for the time covered by population projections (fig.
@@
 Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset.
-The deterministic model keeps the 3% limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
+The deterministic model keeps the 3%-limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.
 Figure 6: Illustration of generational inequality, deterministic model: Lifetime-average cost relative to GDP by year of birth in the different model specifications up to 2200.
@@
 Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP.
-This is because the calibration does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change.
+This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change.
 Absolute cost and its distribution over time are irrelevant to the objective function.
 This effect is driven by the aggregation of discounted welfare into intergenerational welfare.
-Considering stochastic interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost.
-This indicates that a repeated re-calibration of the DICE model to current interest rate levels will amplify intergenerational inequality.
+Considering uncertainty in the rate of time-preference and thus overall interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost.
+This indicates that a repeated re-calculation of the DICE model to current interest rate levels will amplify intergenerational inequality.
 We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost.
-Funding of abatement cost may be considered as modelling the backing of abatement costs by government loans.
+Funding of abatement cost may be considered as modeling the backing of abatement costs by government loans.
 We find that this may have a positive effect on intergenerational equity in terms of the cost distribution as it improves the alignment of abatement and damage cost.
-The non-linear discounting of damage cost may be considered as modelling financing risk of damage costs, where very large cost have higher financing cost.
+The non-linear discounting of damage cost may be considered as modeling financing risk of damage costs, where very large cost have higher financing cost.
 The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
@@
 First, the representation of damages could be improved [47, 11, 12].
-Second, the lack of regional resolution compared to, e.g., econometric damage estimates [50, 51] ignores regional inequalities in exposure and potential distributional effects of projected costs [52].
+Second, the lack of regional resolution compared to, e.g., econometric damage estimates [48, 49] ignores regional inequalities in interest rate dynamics, exposure, and potential distributional effects of projected costs [50].
 Since our study aims to introduce methods such as limiting total costs of climate change by a proportion of GDP to the debate on mitigation pathways, these limitations could be addressed by including the presented concepts in more complex assessments of mitigation pathways [31, 51].
-Modifying the DICE model in a modular way, we address the issue that the classical model does not consider equity between generations.
-The proposed extensions accounting for the funding cost of abatement and increased financing cost of large damages can improve this aspect of the intertemporal optimization.
-This is amplified when considering stochastic interest rate risk.
-Constraining the total costs of mitigation and damages relative to GDP could be used as a simple heuristic to ensure intergenerational equity of climate mitigation.
+Modifying the DICE model in a modular way, we address the issue that the classical model does not consider uncertainty in interest rates and its effects on equity between generations.
+The presented extensions accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality.
+This is amplified when considering stochastic interest rate risk due to convexity induced by the DICE objective function.
+Constraining the total costs of mitigation and damages relative to GDP could be used as a simple objective to ensure intergenerational equity of climate mitigation.
 4 Methods
-The starting point of our investigations and extensions is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
+The starting point of our study is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
 Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model.
-For the equi-distant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered.
+For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered.
 Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$.
@@
 abatement cost and damage cost model) as in DICE-2016.
-A detailed description of our implementation of the DICE model (and its extensions) can be found in [39].
-We use the same notation, except that we denote the cost functions with $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
-We will recall the calibration procedure first in Section 4.1, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
-We then present metrics to analyse the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.2, since the question at which time cost occur is a question of inter-generational equity.
-We conclude presenting different extensions to the model, which change the distribution of cost.
-We discuss these in Section 4.3.
+A detailed description of our implementation of the DICE model and its extensions can be found in [40].
+We use the same notation, except that we denote the cost functions by $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
+We will first recapitulate the role of the time-preference rate $\rho$ in the DICE model in Section 4.1 as a background to our extension with stochastic time-preference rates.
+Second, we recall the optimization procedure in Section 4.2, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
+Third, we present metrics to analyze the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.3.
+Following upon this concepts, we present our extensions to the model, which change the distribution of cost in Section 4.4.
 We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
-4.1 Model Calibration
+Finally, we introduce the concept of convexity from mathematical finance in section 4.5.
+4.1 Time Preference Rate
+The DICE model extends a standard neoclassical optimal growth model (the Ramsey model) with economic impacts from climate change.
+4.1.1 Recapitulation: Time Preference Rates and the Ramsey Equation
+Given that the utility $U$ is a function of the time $t$-consumption $C(t)$, the discounted utility is
+$U(C(t))\cdot\exp(-\rho t)\text{.}$
+If we assume a declining marginal utility
+$\mathrm{d}U(C))\ =\ -\eta\cdot U(C)\cdot\mathrm{d}C\text{,}$
+where $\eta$ denotes the elasticity of marginal utility, then we have
+$\frac{\mathrm{d}}{\mathrm{d}t}\left(U(C(t))\cdot\exp(-\rho t)\right)$
+$\ =\ -\eta\cdot U(C(t))\cdot\frac{\mathrm{d}C(t)}{\mathrm{d}t}\cdot\exp(-\rho t)-\rho\cdot U(C(t))\cdot\exp(-\rho t)$
+$\ =\ -(\eta\cdot g+\rho)\cdot U(C(t))\cdot\exp(-\rho t)\ =\ -r\cdot U(C(t))\cdot\exp(-\rho t)\text{.}$
+with
+$r\ =\ \eta\cdot g+\rho\text{,}$ (1)
+where $g=\frac{\mathrm{d}C(t)}{\mathrm{d}t}$ In this context, $r$ is called the social discount rate and Equation (1) is the Ramsey equation.
+The (optimal) growth of the consumption $g$ depends on the optimal emission and savings paths, and is an outcome of the model optimization.
+The elasticity of marginal utility $\eta$ is an input to the model and usually empirically estimated [52].
+For the link between time preference and social discount rate see also [12].
+The time-preference $\rho$ is an exogenous parameter and subject to controversial discussions.
+That $\rho$ is perceived as exogenous is already transparent from Nordhaus 2017 paper on the DICE model [7], showing a dependency of the social cost of carbon on the chosen (growth-adjusted) discount rate.
+In [53] an analysis is performed on the dependency of the social cost of carbon on a distribution of the discount rate. Here, the Bond market interest rate is used as an input to the model.
+While [53] considers a distribution of discount rates derived from Bond market rates, there is no feedback effect between the integrated assessment model and the given interest rates.
+The novelty of our extensions is to consider a stochastic process of future uncertain discount rates and consider its interaction with an integrated assessment model.
+This models the possibility that policymakers adapt the abatement strategy to changes in interest rates.
+Such an adaptation is plausible, because already the static analysis shows that the optimal abatement strategy depends on the level of the interest rate.
+We model the time-preference $\rho$ by a classical interest rate model, the Hull-White model.
+This can be seen as a model for fluctuations in the time-preference, e.g., induced by fluctuation in the financial market.
+Details are presented in section 4.4.1
+4.2 Model Objective
+The objective of the DICE model is to find the abatement rate $t\mapsto\mu(t)$ and the savings rate $t\mapsto s(t)$ that maximize a given objective function.
 The classical objective function of the DICE model is the integrated discounted welfare
@@
 In the following we will consider an extension of the model where $N$ and $V$ are stochastic processes. This requires that we re-define the objective function as (2) is a random variable.
-For the calibration we consider the expectation operator
+For the optimization we consider the expectation operator
 $\mathrm{E}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{,}$ (3)
@@
 4.2.2 Abatement Policy - Classical DICE Model
-The model calibration determines the optimal abatement policy $t\mapsto\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (1), (2).
-In the time-discrete classical DICE model the calibration determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t)\right)$.
+The model optimization determines the optimal abatement policy $\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (2), (3).
+In the time-discrete classical DICE model the optimization determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t_{i})\right)$.
 For a model time horizon of $T=500\leavevmode\nobreak\ \text{years}$ and a time-discretization step of $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$ this results in 499 free parameters for the abatement policy (and another corresponding set for the savings rate).
@@
 4.3 Cost
-To characterize the resulting optimal mitigation paths, we utilise derived quantities.
+To characterize the resulting optimal mitigation paths, we utilize derived quantities.
 A prominent example of such a derived quantity is the social cost of carbon, $SCC(t)$.
@@
 4.3.1 Distribution of Cost
-We analyse the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
+We analyze the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
 $C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
-To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the calibration’s objective function.
+To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the optimization’s objective function.
 4.3.2 Sensitivity of Damage and Abatement Cost to Policy Changes
@@
 The product of both weights represents the cost-to-value-weight.
-The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfils
+The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfills
 […]
 From eq.
-6 we see why the model calibration may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered.
+7 we see why the model optimization may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered.
 Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq.
@@
 This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$.
-As the model calibration is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity.
+As the model optimization is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity.
 The minus in (9) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
 $\rho_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (9)
-From Equation 5 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density.
+From Equation 6 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density (Figure 8).
 Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
 Figure 8: The density $t\mapsto\rho_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
-In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement (Figure 8), indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
+In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement.
+This can be observed in Figure 2, indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
 Sensitivity of Cost to Time of Reaching 100 % Abatement
@@
 2.
-4.2.3 Cost per GDP
-To measure the burden born by a generation, we consider the relative total damage and abatement cost per GDP.
-Putting nominal values relative to the GDP is a common measure, e.g., to assess national debt [45].
-Here, it reflects that a more wealthy generation can carry a greater burden, mimicking an effort sharing scheme and also considering the greater capacity of future generations under the expected baseline growth in DICE.
-In our numerical experiments we measure the instantaneous cost-per-GDP $C(s)/GDP(s)$ as well as the lifetime-average cost-per-GDP $\overline{C}(s)/\overline{GDP}(s)$.
-To allow for a netting inside a generation one may consider the (discounted) average value over a generation lifetime.
-For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation.
-We then consider the lifetime average cost and GDP.
-$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
 4.4 Model Extensions
@@
 4.4.1 Stochastic Interest Rates
-The classical way in which interest rates are modelled in an IAM is via a discount factor.
+The classical way in which time-preference rates are modeled in an IAM is via a discount factor.
 A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
-$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
-We model stochastic interest rates $r$.
+$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
+We model stochastic interest rates (time-preference rates) $rho$.
 Our implementation allows to use a general discrete forward rate model (LIBOR Market Model).
@@
 Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates).
-For exemplary parameterization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$.
+For exemplary parametrization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$.
 While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively.
@@
 If interest rates are stochastic but the function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM since
-$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (10)
-with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}r(s)\mathrm{d}s\right)\right)\right)$.
+$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (11)
+with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\right)$.
 Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics.
 This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process.
-[^2] Our choice to make interest rates stochastic is only exemplary.
+[^2] Our choice to make interest rates stochastic through uncertain time preference rates is only exemplary.
 The interest rate is an important factor in linking present abatement cost to the avoided future damage cost.
 While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
-We use a classical Hull-White model, that is
-$\begin{split}\mathrm{d}r(t)&\ =\ (\theta(t)-a(t)r(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad r(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ r(t)\ \mathrm{d}t\text{,}\end{split}$ (11)
+To model the uncertainty of the rate of time-preference $\rho(t)$, we use a classical Hull-White model [39, 55], that is
+$\begin{split}\mathrm{d}\rho(t)&\ =\ (\theta(t)-a(t)\rho(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad\rho(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ \rho(t)\ \mathrm{d}t\text{,}\end{split}$ (12)
 where an exact time-discretization scheme [56] was used.
-The model’s mean-reversion level $\theta$ is such that the numéraire-relative zero coupon bond prices are martingales and their expectation agrees with the classical non-stochastic model.
-In other words, our stochastic interest rate model is such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
+Our stochastic interest rate model is parameterized such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
 4.4.2 Stochastic Abatement Policy
-Assuming that a change in the interest rates triggers a re-calibration of the model, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic.
+Assuming that a change in the interest rates triggers a re-calculation of the model optimal abatement policy, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic.
 The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates.
@@
 An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization.
-As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $r(t)$.
-A simple example of a stochastic abatement model is a parametric one, where the abatement speed is a (linear) function of the interest rate level $r$, e.g.,
-$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot r(t,\omega)\right)\cdot t,1.0\right)\text{.}$ (12)
-The simplicity of the models allows us to interpret the parameters.
+As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $\rho(t)$.
+To obtain a robust optimization, reducing the risk of over-fitting or numerical instabilities, we propose a parametric stochastic abatement model. Here, we use the model
+$\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot\rho(t,\omega)\right)\cdot t,1.0\right)\text{,}$ (13)
+where the abatement speed is a (linear) function of the interest rate level $\rho$.
+The model can be seen as a first-order approximation of general policies, where the abatement is allowed to depend in the level of the discount rate $\rho$.
+The simplicity of the model allows us to interpret the parameters.
 We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
 $a_{0}\ =\ \frac{\partial\mu(t)}{\partial t}\big{|}_{r=0}\text{,}\qquad a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
-Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (4) simply due to the additional degree of freedom $a_{1}$ in the abatement optimisation.
-The calibration of the optimal model (12) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
+Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (5) simply due to the additional degree of freedom $a_{1}$ in the abatement optimization.
+The optimization of the model (13) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
 4.4.3 Funding of Abatement Costs
@@
 We introduce the option of funding of abatement cost.
-As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest rate corresponds to the current discount rate.
+As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest corresponds to the current discount rate.
 Our model allows to apply a funding spread, i.
@@
 higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon.
-In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, as introduced in DICE-2023 [38].
+In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, comparable to the modifications in DICE-2023 [38].
 We define $C_{\mu}(t)$ as the instantaneous abatement costs in time $t$, i.e., the quantity that was formerly denoted by $C_{\mathrm{A}}(t)$.
@@
 For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, we recover the classical model with $C_{\mu}=C_{\mathrm{A}}$.
-In a standard model for risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting.
-However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated and due to the time-preference included in the utility function.
+In a standard risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting.
+However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated as well as due to the time-preference included in the utility function.
 Since damages occur instantaneously, we do not consider funding of these. The total cost is given by $C(t):=C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)$.
@@
 Thus additional cost may occur to secure the unsecured funding [42].
-Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger cost.
+Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger damages.
 We model this by optionally adding non-linear financing cost using a non-linear funding model [42].
@@
 Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions.
-For our application, this means that larger cost get a larger weight.
+For our application, this means that larger damages get a larger weight.
 Our model modification is now a modification of the damage cost.
@@
 As the abatement cost are usually smaller and part of a more planned process, we do not consider a default compensation factor for the abatement cost.
+4.4.5 Cost per GDP
+To measure the burden born by a generation, we consider the relative total damage and abatement cost per GDP.
+Putting nominal values relative to the GDP is a common measure, e.g., to assess national debt [57].
+Here, it reflects that a more wealthy generation can carry a greater burden, mimicking an effort sharing scheme and also considering the greater capacity of future generations under the expected baseline growth in DICE.
+In our numerical experiments we measure the instantaneous cost-per-GDP $C(s)/GDP(s)$ as well as the lifetime-average cost-per-GDP $\overline{C}(s)/\overline{GDP}(s)$.
+To allow for a netting within a generation one may consider the (discounted) average value over a generation lifetime.
+For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation.
+We then consider the lifetime average cost and GDP.
+$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
+4.5 Convexity
+Our numerical analysis shows that a continuous adaptation of the abatement strategy may increase the risk and expectation of intergenerational inequality measures as cost-per-GDP.
+The expected emissions/damage/cost are higher than the emissions/damage/cost calculated for the expected interest rate level.
+Such an effect may come as a surprise.
+However, it is well known from mathematical finance, known under the name convexity.
+The effect corresponds to Jensen’s inequality, that the expectation of a convex function of a random variable is higher than the convex function applies to the expectation of the random variable.
+Here, the random variable is the exogenous given discount rate, or stochastic process of the interest rates, and the integrated assessment model acts as a convex function.
+In other words: DICE has convexity in the interest rates once the abatement strategy depends on the interest rate level.
 Supplementary Material
@@
 Acknowledgements
-We thank Max Singhoff for exploring the calibration of the DICE model with the ADAM solver used in our numerical experiments.
+We thank Max Singhoff for exploring the optimization of the DICE model with the ADAM solver used in our numerical experiments.
 L.Q.
@@
 figureSupplementary figures
+Variation of the Interest Rate Volatility
+Supplementary Figure 1: Stochastic DICE Model with Different Interest Rate Volatilities ($\sigma=0.2\%$ (red), $\sigma=0.3\%$ (blue), $\sigma=0.4\%$ (green)).
+Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost).
+Shaded area shows the 10th to 90th percentiles.
+Variation of the Interest Rate Level
 Supplementary Figure 2: Increasing interest rate level increases the time of 100% abatement $T^{\mu=1}$.

### v4

@@
-Positive Effects of Financing
+Intergenerational Equitable Climate Change Mitigation:
 Abstract
-Today’s decisions on climate change mitigation affect the damage that future generations will bear.
-Discounting future benefits and costs of climate change mitigation is one of the most critical components of assessing efficient climate mitigation pathways.
-We extend the DICE model with stochastic discount rates to reflect the uncertain nature of discount rates.
-Stochastic rates give rise to a stochastic mitigation strategy, resulting in all model quantities becoming stochastic.
-We show that the optimization procedure of the DICE model induces intergenerational inequality: lacking a mechanism to regulate burden, future generations have to bear higher costs from abatement and damage relative to GDP.
-Further, we show that considering uncertainty of discount rates and their feedback to abatement policies, which can be interpreted as successive re-calculation, increases intergenerational inequality (and adds additional risks).
-Motivated by this, we consider additional financing risks by investigating two modifications of DICE.
-We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing.
-To conclude our discussion of options to improve intergenerational equity in an IAM, we propose a modified optimization to keep costs below 3% of GDP, resulting in more equal distribution of efforts between generations.
+Climate mitigation decisions today affect future generations, raising questions of intergenerational equity.
+Integrated assessment models (IAMs) rely on discounting to evaluate long-term policy costs and benefits.
+Using the DICE model, we quantify how optimal pathways distribute abatement and damage costs across cohorts.
+Unconstrained optimization creates intergenerational inequality, with future generations bearing higher costs relative to GDP.
+Extending the model with stochastic discount rates, we show that discount-rate uncertainty significantly amplifies this inequality.
+We consider two independent extensions: the financing of abatement costs and the modeling of nonlinear financing costs under large damages.
+Both extensions can materially improve intergenerational equity by distributing mitigation efforts more evenly.
+As an illustration, we present a modified DICE model whose optimal pathway limits generational costs to 3 % of GDP, leading to more equitable effort sharing.
+Our proposed model extensions are model-agnostic, applicable across IAMs, and compatible with alternative intergenerational equity metrics.
+[…]
+$\printcontents[main]$l1
 1 Introduction
-Climate change is one of the most significant risks of the next centuries [1], as demonstrated by its manifold impacts on society [2].
-Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide [3], a transition to (net) zero emissions is necessary to limit global warming to any constant level.
-A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world.
-One well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
-We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12].
+Climate change is one of the largest risks of the next centuries [1], as demonstrated by its manifold impacts on society [2].
+Owing to the persistent nature of the main greenhouse gas carbon dioxide [3] on human time scales, a transition to (net) zero emissions is necessary to limit global warming to any constant level.
+A well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of climate and economic systems [4, 5].
+We focus on the DICE model [6, 7, 8] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [9, 10, 11, 12, 13].
 While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [14, 15, 16, 17, 18, 19, 20], natural feedback processes such as permafrost [10] or abstract catastrophic risks [21, 22].
-Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimization function have a strong influence on the resulting optimal pathways [22, 23, 24, 25, 26].
+Because most of the damages from climate change will occur over long time scales, value-based decisions about the discount rate and the optimization function have a strong influence on the resulting optimal paths [23, 24, 25, 26, 27].
 Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [14] or inequalities in the distribution of damages [28].
-Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation.
-Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels.
 Discount rates, in particular, are an essential component of IAM modeling.
-Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues.
+Thus, the discussion between proponents of a descriptive approach based on observation of market returns as in DICE [7] versus a normative approach to discounting as in the Stern review [23] continues.
 The assumptions on discounting have already been discussed in the context of intergenerational [29] and intragenerational inequality [30].
@@
 These uncertainties are translated into certainty-equivalent discount rate, such that the overall model remains deterministic.
-As we show in eq.
-11, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
-Complementary to the potential uncertainties in future consumption growth and riskiness of climate investments, we consider uncertainty about the rate of time preference $\rho$ itself.
-For this, we use a standard interest-rate model - the Hull-White model [39] to model stochastic fluctuation of the rate of time preference $\rho$, interacting with an abatement policy adjusting to this uncertainty.
-1.1 Extensions to the DICE Model
+As we show in eq.˜9, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
+In contrast to this deterministic treatment of interest rate risk, we consider uncertainty about the rate of time preference $\rho$ by modeling it as a stochastic process using a standard interest rate model.
+Introducing adaptation of the abatement policy, all quantities become stochastic.
+We then show that the optimal mitigation pathway derived from an IAM may exhibit substantial intergenerational cost imbalances, which are critically exacerbated by uncertainties in interest rates.
+We propose two novel model extensions that markedly improve cost-based intergenerational equity, thus enhancing the robustness and fairness of modeling global climate mitigation, providing a crucial political impulse for effective climate action.
+fig.˜1 provides an overview of the extensions.
+Here, we interpret intergenerational equity as ensuring that no cohort bears a disproportionate share of climate-related mitigation and damage costs relative to its economic capacity.
+To allow meaningful comparisons between generations of differing wealth and productivity, we measure costs as a percentage of GDP, following principles in public finance for sustainable debt and investment burdens [35].
+As an illustrative benchmark rather than a normative prescription, we adopt a limit of 3 % of GDP for combined mitigation and damage costs, drawing on similar proposals for a “fair share” of national income devoted to climate action and its consistency with typical upper bounds for long-term public investment commitments.
+We recognise that alternative justice principles, such as egalitarian, prioritarian, or sufficiency-based approaches [29], would lead to different equity metrics and thresholds.
+Our results are therefore one contribution to the broader debate on embedding intergenerational fairness in integrated assessment models, but the model extensions we propose can be readily combined with such alternative metrics.
+{delete}
+2 Method: Extensions to the DICE Model
 We examine the optimization of the DICE model with respect to uncertainty in the rate of time-preference, and its consequences for intergenerational equity.
-Measuring the costs per GDP (section 4.4.5) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations.
+Measuring the costs per GDP (section˜5.2.2) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations.
 We use an implementation of the DICE model with a flexible time discretization [46] and couple this model with a comprehensive modeling toolbox for financial markets [36] to enable the integration of valuation methods for complex financial products in IAMs.
@@
 As abatement is a planned investment process, we assume that these costs may be financed by loans with uncertain interest rates.
-Furthermore, as large projects and investments often incur additional costs [43, 44], we introduce increasing financing costs for large payments via a non-linear discounting.
+Furthermore, as large projects and investments often incur additional costs [hallegatte_why_2007, shane_construction_2009], we introduce increasing financing costs for large payments via a non-linear discounting.
 These higher costs for high damages can be viewed as (additional) financing risks.
-We provide an overview of the extensions in fig.
-1.
-More background is provided in the methods section.
 Basic model DICE-2016-R [7]  –  re-implementation with improvements
@@
 3.1 Intergenerational Inequality
-To find the optimal mitigation pathway, the standard DICE model equates the increment of the cost of damage $\mathrm{d}C_{\mathrm{D}}(t)$ to the increment of the cost of abatement $\mathrm{d}C_{\mathrm{A}}(t)$, both weighted by the welfare-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor $\frac{N(0)}{N(t)}$,
-$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (8)
-Considering a simplified 1-parameter abatement model parameterizing the abatement policy by $T^{\mu=1}$, the time by which 100% abatement is achieved, this simplifies to
-$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (10)
-The conditions (8) and (10) are plausible if a single agent optimizes its cost but may be problematic if the costs are distributed between different generations.
-Note that the two individual contributions in eq.
-10 occur at different times.
-fig.
-2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
+The optimality condition of the IAM’s emission path balances the sensitivity of the two cost components, abatement cost and damage cost, to policy changes (see eq.˜5).
+We individually depict their temporal distribution: the weighted marginal change in abatement and damage cost with respect to a policy change in fig.˜2 and find that an unconstrained IAM determines the optimality of the emission path by balancing the near-time abatement cost changes with the future time damage cost changes.
 Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$.
-The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
-Hence, marginal gains of one generation are equated to marginal losses of another generation.
-The optimization does not equalize the burden but the weighted marginal burden.
-In this sense, the optimization is indifferent to the temporal distribution of the absolute burden.
-This may generate intergenerational inequality.
+The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost.
+The two sensitivities come with a significant time shift.
+This may already constitute intergenerational inequality, as within that optimization the absolute burden of near-time abatement cost may be much lower than future-time damage cost.
+As changes in the abatement policy at time $s$ affect the abatement cost at time $s$ but damage cost at all future times $t\geq s$, we define $\gamma_{\mathrm{D}}(s;t)$ as the damage cost in $t$ per abatement cost in $s$ induced by a change in the abatement policy $\mu(s)$, see (6).
+We show that at the optimal emission path $\gamma_{\mathrm{D}}(s;t)$ is a probability distribution; a proof can be found in Supplement Section B.
+This allows us to define the expected time-delay between an abatement and its effect on future damage cost, and we are able to determine it numerically.
+fig.˜3 depicts the density $\gamma_{\mathrm{D}}(s)$.
+Figure 3: The density $t\mapsto\gamma_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
+We find that in the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement.
+This observation agrees with fig.˜2, indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
+We depict the temporal distribution of the total cost in fig.˜4, fig.˜5.
+3.2 Cost per GDP
+Since the optimal emission path balances marginal cost changes only but does not consider cost levels, we like to analyze the temporal distribution of the cost.
+To measure the burden borne by a generation, we consider the total cost per GDP.
+Putting nominal values relative to the GDP is a common measure, for example, to assess national debt [35].
+Here, it reflects that a wealthier or more productive generation can carry a greater burden, mimicking an effort-sharing scheme, and also considering the greater capacity of future generations under the expected baseline growth in DICE.
 3.3 Including stochastic interest rates into DICE
-We see in Equation 8 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the discount factor $\frac{N(0)}{N(t)}$ (i.e., the numeraire ratio) are important to determine the inter-temporal optimal pathway.
-Including a stochastic interest rate model, the numéraire $N$ becomes stochastic.
-Allowing for (stochastic) adaptation of the abatement policy $\mu$ to the interest rate level, all model quantities become stochastic.
-We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$ details are provided in the methods (12).
-The model generates the same expected discount rate as for the deterministic DICE model.
-This implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model.
-As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [40].
-However, allowing for a stochastic abatement policy that adapts to the interest rate level, e.g., under the model (13) presented in the methods, we see an adverse effect: the expected abatement is smaller, the expected emissions are higher and the expected total cost increase significantly for later years.
-Note that restricting the stochastic abatement model (13) to the case $a_{1}=0$ agrees with the abatement model (5).
-The effect is hence solely created by allowing the change of abatement speed by interest rate change $a_{1}$ to be a free parameter.
-fig.
-3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage and total cost-per-GDP.
-These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
-Figure 3: Classical DICE (grey), Classical DICE with reduce Abatement Model (5) (blue) and Stochastic DICE Model (green).
+Discounting plays a fundamental role in the linkage of present and future cost. We assess how uncertainty in discount rates alters the distribution of cost in the optimal emission paths.
+We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$.
+If interest rates are stochastic but the utility function remains deterministic, adding stochastic interest rates does not change the interaction with the IAM, and hence the optimal emission path.
+However, this changes if one assumes that the abatement policy is adapted to the (stochastic) changes in the interest rate level.
+We allow for a stochastic abatement policy that adapts to the interest rate level, using the parametric model (10).
+This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
+With the stochastic abatement policy, we see an adverse effect: the expected abatement is smaller, the expected emissions are higher, and the expected total cost increases significantly for later years.
+Since by restricting a parameter of the stochastic abatement policy ($a_{1}=0$ in (10)) it agrees with the deterministic abatement model (2), it is apparent that the effect is created solely by allowing the abatement speed to depend on the interest rate level (i.e., $a_{1}$ being a free parameter in (10)).
+fig.˜4 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage, and total cost-per-GDP.
+Figure 4: Classical DICE (grey), Classical DICE with reduced Abatement Model (2) (blue) and Stochastic DICE Model (red).
 Restricting the abatement policy to the 1-parameter model (2) has a comparably small effect on the emission pathways.
@@
 Shaded area shows the 10th to 90th percentiles.
+The deterministic model (blue) uses the certainty-equivalent discount rate $\bar{r}$ from (9); therefore, differences in the red line arise solely from interest-rate–driven policy adaptation ($a_{1}\neq 0$) and damage nonlinearity, not from a change in average discounting.
 This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles).
-This is qualitatively independent of the interest rate level.
-Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (11) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates.
-Note that for a convex function $f$ and a random variable $X$ Jensen’s inequality implies $f(\mathrm{E}(X))\leq\mathrm{E}(f(X))$.
-The strength of the effect increases with increasing interest rate volatility.
-A corresponding analysis for different volatility levels can be found in Variation of the Interest Rate Volatility of the supplementary information.
-For the limits cast of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
-2.3 Including funding of abatement and funding risks of damage into DICE
-As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future.
-We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time, resulting in interest that has to be paid.
-Second, we include a non-linear dependency of the discount factor for the damage cost on the magnitude of the cost.
-This can be viewed as the additional cost associated with financing risk of very large amounts [42].
-Details on this extension and parameters are provided in the methods section 4.4.4.
-We find that both extensions improve intergenerational equity in the classical deterministic DICE model, fig.
-4, and in the DICE model with stochastic interest rate, fig.
-5.
-In the latter, intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
-Figure 4: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
-Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
-Shaded area shows the 10th to 90th percentiles (VaR).
-Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs.
-Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$).
+In other words, the DICE model exhibits convexity with respect to the interest rates.
+This effect is qualitatively independent of the interest rate level.
+The strength of the effect increases with an increase in interest rate volatility, LABEL:fig:pathways_stochastic_variation.
+For the limit case of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
+3.4 Improving Intergenerational Equity
+We investigate the effect of the two model extensions “funding of abatement cost” (Section 5.3.3) and “non-linear discounting of damage cost” (Section 5.3.4) on the distribution of cost in the deterministic and stochastic model, i.e., how they improve cost-related intergenerational equity.
+The parameter choices are exemplary; other experiments with similar results can be reproduced in the published source code.
+3.4.1 Funding of Abatement
+We allow funding of abatement costs to reduce the time lag between abatement cost and damage cost.
+As we account for interest for the funding, this does not alter the effect of discounting but brings the two components to similar economical regions.
+Shifting cost via a loan to the future may be considered adverse to intergenerational equity; however, our numerical analysis indicates that it is beneficial as it incentivizes higher abatement.
+In a numerical experiment, with a funding period of 60 years with equal-value annual repayment (including interest), funding of abatement reduces the cost in the very far future without substantial increase in the near future (fig.˜5, left).
+An even more notable contribution is that it significantly reduces the cost risk if discount rates are stochastic (fig.˜5, right).
+However, with respect to intergenerational equity, improvements are somewhat unbalanced as a funding of abatement still leaves a high burden for near future generations.
+3.4.2 Funding Risks of Damage
+We add a non-linear discounting of damages.
+This model extension allows the discount factor to depend on the magnitude of the cash flow, mimicking that funding of larger cash flows requires a larger premium to compensate for default risk or other frictions, a default compensator.
+As such, larger damages get a larger weight; consequently, the abatement policy optimization will prefer emissions paths that avoid very large damages, which is beneficial with respect to intergenerational equity.
+In our numerical experiment, we consider a significant default compensator for damage cost over 3% of the GDP, effectively limiting cost of optimal emission paths to $3\%$.
+We then find strong improvements to intergenerational equity in the deterministic DICE model and the DICE model with a stochastic interest rate, as shown in fig.˜5.
+Intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
+The associated model quantities (emission, abatement and damage) are depicted in LABEL:fig:pathways_deterministic_with_extension and LABEL:fig:pathways_stochastic_with_extension.
+Figure 5: The cost-per-GDP in a the model with deterministic interest rates and b the model with stochastic interest rates, for the otherwise unmodified model (red) and the model with the extension “funding of abatement cost” (blue) and “GDP limited non-linear discounting” (green).
+Shaded area shows the 10th to 90th percentiles.
+Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, the present generations still face the lowest costs.
 For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [37] to calculate burdens experienced during each birth-year cohorts projected lifetime.
-We measure the burden in percentage-of-GDP (section 4.4.5), where we choose 3% as the neutral level (green), [46].
-While differences are small for the time covered by population projections (fig.
-6 and fig.
-7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century.
+We measure the burden in percentage-of-GDP (section˜5.2.2), where we choose 3% as the neutral level (green), [38].
+While differences are small for the time covered by population projections (fig.˜6 and fig.˜7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century.
 Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset.
@@
 4 Discussion
-We analyze the DICE integrated assessment model and derive metrics to assess intergenerational equity of the optimal emission pathways calibrated by the model.
-We extend the DICE integrated assessment model and compare several extensions for their effect on intergenerational equity.
-Our analyses apply to integrated assessment models in general and could also be extended to benefit-cost analysis of climate mitigation.
-We provide a complete open-source implementation of the model illustrating all our extensions, utilizing Monte-Carlo implementation of standard interest rate models for simulation and (stochastic) algorithmic differentiation to efficiently calculate sensitivities of model quantities against each other.
-Our first observation is that the DICE model can significantly increase the policy cost measured in terms of cost-per-GDP.
-This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change.
-Absolute cost and its distribution over time are irrelevant to the objective function.
-This effect is driven by the aggregation of discounted welfare into intergenerational welfare.
+We analyzed the DICE integrated assessment model and derived metrics to assess the temporal cost structure and thus intergenerational equity of the optimal emission pathway.
+We extend the DICE integrated assessment model and compare several extensions for their effects on intergenerational equity.
+Our methods and extensions apply to integrated assessment models in general and can also be extended to the benefit-cost analysis of climate mitigation.
+Our first observation is that the DICE model optimization results in an unequal temporal distribution of policy costs measured in terms of cost-per-GDP.
+This is because the optimization does not consider any (cost-related) intergenerational equity but balances the marginal aggregated cost changes induced by a policy change, while absolute cost and its distribution over time are irrelevant to the objective function.
 Considering uncertainty in the rate of time-preference and thus overall interest rates, this effect is amplified when the stochastic adaptation of the abatement strategy generates a significant risk of increased cost.
-This indicates that a repeated re-calculation of the DICE model to current interest rate levels will amplify intergenerational inequality.
+This indicates that repeated re-calculation of the DICE model to current interest rate levels amplifies intergenerational inequality.
 We consider natural model extensions by allowing the funding of abatement cost and non-linear discounting of damage cost.
@@
 The non-linear discounting of damage cost may be considered as modeling financing risk of damage costs, where very large cost have higher financing cost.
-The non-linear damage cost can be used as a constrain that allows to effectively limit the generational cost to a specific percentage of the GDP.
+The non-linear damage cost can be used as a constraint that allows to effectively limit the generational cost to a specific percentage of the GDP.
 Using a simple IAM includes several caveats.
@@
 Modifying the DICE model in a modular way, we address the issue that the classical model does not consider uncertainty in interest rates and its effects on equity between generations.
-The presented extensions accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality.
+The extensions presented accounting for the funding cost of abatement and increased financing cost of large damages reduce intergenerational inequality.
 This is amplified when considering stochastic interest rate risk due to convexity induced by the DICE objective function.
 Constraining the total costs of mitigation and damages relative to GDP could be used as a simple objective to ensure intergenerational equity of climate mitigation.
+Our results can also be interpreted as supporting the design of financing mechanisms such as a climate transformation fund that channels dedicated, long-term resources into mitigation and adaptation, see also [45], which builds on the modeling framework developed in the present work and verifies that a climate transformation fund can be funded with less than 3 % of the GDP.
+Implementing these mechanisms in practice would require collaboration between policymakers, financial institutions, and affected communities to ensure that they are economically sustainable, politically viable, and perceived as fair between generations.
 5 Methods
-The starting point of our study is the 2016 implementation of the classical DICE model [6, 7], since the DICE-2023 description [38] was only published parallel to this research.
-Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is re-defined as being an Euler-discretisation of a time-continuous model.
-For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered.
-Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$.
-The differences to the classical DICE-2016 are negligible for our work.
-Except where otherwise stated, we use the same model components (e.g.
-abatement cost and damage cost model) as in DICE-2016.
+The starting point of our modeling exercise is DICE-2016 [6, 7].
+Except where otherwise stated, we use the same model components as in DICE-2016.
 A detailed description of our implementation of the DICE model and its extensions can be found in [46].
-We use the same notation, except that we denote the cost functions by $C$ [^1] In [7] the letter $C$ was used for consumption, which we do not reference here.
-We will first recapitulate the role of the time-preference rate $\rho$ in the DICE model in Section 4.1 as a background to our extension with stochastic time-preference rates.
-Second, we recall the optimization procedure in Section 4.2, where we introduce a parameter-reduced form of the abatement policy to use for our analysis.
-Third, we present metrics to analyze the temporal distribution of cost and the temporal distribution of cost-sensitivity to policy changes in the classical DICE model in Section 4.3.
-Following upon this concepts, we present our extensions to the model, which change the distribution of cost in Section 4.4.
-We perform our analysis for the classical (deterministic) DICE model, as well as a model extended with stochastic interest rates and a stochastic abatement policy.
-Finally, we introduce the concept of convexity from mathematical finance in section 4.5.
-4.1 Time Preference Rate
-The DICE model extends a standard neoclassical optimal growth model (the Ramsey model) with economic impacts from climate change.
-4.1.1 Recapitulation: Time Preference Rates and the Ramsey Equation
-Given that the utility $U$ is a function of the time $t$-consumption $C(t)$, the discounted utility is
-$U(C(t))\cdot\exp(-\rho t)\text{.}$
-If we assume a declining marginal utility
-$\mathrm{d}U(C))\ =\ -\eta\cdot U(C)\cdot\mathrm{d}C\text{,}$
-where $\eta$ denotes the elasticity of marginal utility, then we have
-$\frac{\mathrm{d}}{\mathrm{d}t}\left(U(C(t))\cdot\exp(-\rho t)\right)$
-$\ =\ -\eta\cdot U(C(t))\cdot\frac{\mathrm{d}C(t)}{\mathrm{d}t}\cdot\exp(-\rho t)-\rho\cdot U(C(t))\cdot\exp(-\rho t)$
-$\ =\ -(\eta\cdot g+\rho)\cdot U(C(t))\cdot\exp(-\rho t)\ =\ -r\cdot U(C(t))\cdot\exp(-\rho t)\text{.}$
-with
-$r\ =\ \eta\cdot g+\rho\text{,}$ (1)
-where $g=\frac{\mathrm{d}C(t)}{\mathrm{d}t}$ In this context, $r$ is called the social discount rate and Equation (1) is the Ramsey equation.
-The (optimal) growth of the consumption $g$ depends on the optimal emission and savings paths, and is an outcome of the model optimization.
-The elasticity of marginal utility $\eta$ is an input to the model and usually empirically estimated [52].
-For the link between time preference and social discount rate see also [12].
-The time-preference $\rho$ is an exogenous parameter and subject to controversial discussions.
-That $\rho$ is perceived as exogenous is already transparent from Nordhaus 2017 paper on the DICE model [7], showing a dependency of the social cost of carbon on the chosen (growth-adjusted) discount rate.
-In [53] an analysis is performed on the dependency of the social cost of carbon on a distribution of the discount rate. Here, the Bond market interest rate is used as an input to the model.
-While [53] considers a distribution of discount rates derived from Bond market rates, there is no feedback effect between the integrated assessment model and the given interest rates.
-The novelty of our extensions is to consider a stochastic process of future uncertain discount rates and consider its interaction with an integrated assessment model.
-This models the possibility that policymakers adapt the abatement strategy to changes in interest rates.
-Such an adaptation is plausible, because already the static analysis shows that the optimal abatement strategy depends on the level of the interest rate.
-We model the time-preference $\rho$ by a classical interest rate model, the Hull-White model.
-This can be seen as a model for fluctuations in the time-preference, e.g., induced by fluctuation in the financial market.
-Details are presented in section 4.4.1
+We use the same notation, except that we denote the cost functions by CC.
+[^1] In [7] the letter $C$ was used for consumption, which we do not refer to here.
+We redefine the objective function to be suitable for stochastic model extensions; see Section 5.1.
+We then analyze the temporal distribution of cost and the cost sensitivity to policy changes in the classical DICE model in Section 5.2 as an indicator of intergenerational equity.
+Section 5.3 introduces model extensions, which allow to alter the distribution of cost towards intergenerational equitable climate mitigation.
+Most importantly, we allow for a stochastic time preference $\rho$ modeled by a classical interest rate model, as described in section˜5.3.1.
 5.1 Model Objective
-The objective of the DICE model is to find the abatement rate $t\mapsto\mu(t)$ and the savings rate $t\mapsto s(t)$ that maximize a given objective function.
-The classical objective function of the DICE model is the integrated discounted welfare
+The objective of the DICE model is to determine the abatement rate $t\mapsto\mu(t)$ and the savings rate $t\mapsto s(t)$ that maximize a given objective function.
+The objective function of the classical DICE model is the integrated discounted welfare
 $\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\text{.}$ (1)
-Here $N$ denotes the numéraire.
-For the classical model, this means that $N(0)/N(t)$ is the discount factor.
+Here, $V$ denotes the utility and $N$ denotes the numéraire.
+DICE-2016 is recovered for $N(t)=\exp(\rho\cdot t)$.
 We use the more general notion of a numéraire, as we will consider a more general model later.
-DICE-2016 is recovered by $N(t)=\exp(r\cdot t)$.
-4.2.1 Objective Function for the Stochastic Model
-In the following we will consider an extension of the model where $N$ and $V$ are stochastic processes. This requires that we re-define the objective function as (2) is a random variable.
-For the optimization we consider the expectation operator
-$\mathrm{E}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{,}$ (3)
-i.e., the objective function for the stochastic model is the expected welfare.
-For analysis purposes we will plot a risk measure, here the value-at-risk risk, i.e., a percentile of the stochastic welfare,
-$\mathrm{VaR}_{\alpha}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{.}$ (4)
-4.2.2 Abatement Policy - Classical DICE Model
-The model optimization determines the optimal abatement policy $\mu(t)$ and optimal savings rate $s(t)$ that maximizes the integrated discounted welfare (2), (3).
-In the time-discrete classical DICE model the optimization determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t_{i})\right)$.
-For a model time horizon of $T=500\leavevmode\nobreak\ \text{years}$ and a time-discretization step of $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$ this results in 499 free parameters for the abatement policy (and another corresponding set for the savings rate).
-4.2.3 Parametric Abatement Policy - Reduced Model
-For analysis purposes and for comparison with a parametric stochastic abatement model, which we will introduce as part of our model extensions later, we define a parametric abatement model
+For the stochastic model, the objective function will be an expectation of (1) and a quantile for uncertainty analysis.
+5.1.1 Abatement Policy - Classical DICE Model and Reduced Model
+In the time-discrete classical DICE model, optimization determines the optimal parameters $i\mapsto\left(\mu(t_{i}),s(t_{i})\right)$.
+For comparison with a parametric stochastic abatement model, which we introduce later as part of our model extensions, we define a parametric abatement model
 $\mu(t)\ =\ \min\left(\mu(0)+\frac{1-\mu(0)}{T^{\mu=1}}\ t,1.0\right)\text{,}$ (2)
-where the parameter $T^{\mu=1}$ represents the time for reaching 100% abatement.
+where the parameter $T^{\mu=1}$ represents the time required to reach 100% abatement.
 This model has only a single parameter, however it turns out that the optimal abatement policy in the full parameter model is comparably close to a functional form of the type (2).
-4.3 Cost
-To characterize the resulting optimal mitigation paths, we utilize derived quantities.
-A prominent example of such a derived quantity is the social cost of carbon, $SCC(t)$.
-The $SCC$ is the time-$t$ marginal cost of emitting one additional unit of carbon.
-Its unit is $[SCC]=\frac{\mathrm{USD}}{\mathrm{tCO}_{2}}$.
-The “social cost of carbon” is maybe not a suitable measure to access the temporal distribution of the social burden, and hence intergenerational equity, because pricing emissions at the level of the SCC does not cover the cost associate with the climate mitigation paths: the accumulated value $\int_{0}^{T}SCC(t)\cdot E(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$ does not match the accumulated cost $\int_{0}^{T}C(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$, see [54].
-4.3.1 Distribution of Cost
-We analyze the temporal distribution of the cost $C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
-$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{.}$
-To gain a better understanding, we decompose how changes of the abatement policy $\mu$ translate into changes of the cost structure and how these enter into the optimization’s objective function.
+5.2 Metrics Describing Intergenerational Burden
+To characterize the resulting optimal mitigation paths with respect to Intergenerational equity, we consider the total cost.
+5.2.1 Cost
+We analyze the temporal distribution of cost $t\mapsto C(t)$, which is the sum of the abatement cost $C_{\mathrm{A}}$ and the damage cost $C_{\mathrm{D}}$
+$C(t)\ :=\ C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)\text{,}$
+and investigate the interplay of abatement and damage cost.
+5.2.2 Cost per GDP
+We measure the burden borne by a generation by considering the total cost per GDP, as it is common in other contexts, for example, to assess national debt [35].
+Here, it can be considered as mimicking an effort-sharing scheme, that a more wealthy or productive generation can carry a greater burden,
+To allow for a netting within a generation, one may consider the (discounted) average value over a generation lifetime.
+For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation.
+We then consider the lifetime-average cost and GDP,
+$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
+In our numerical experiments, we measure the instantaneous cost-per-GDP $\frac{C(s)}{GDP(s)}$ as well as the lifetime-average cost-per-GDP $\frac{\overline{C}(s)}{\overline{GDP}(s)}$.
 5.2.3 Sensitivity of Damage and Abatement Cost to Policy Changes
-Differentiating the objective function (2) with respect to the abatement policy $\mu$ gives
-$\frac{\mathrm{d}}{\mathrm{d}\mu}\ \int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t$ $\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C(t)}{\mathrm{d}\mu}\ \frac{N(0)}{N(t)}\mathrm{d}t$
-$\ =\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\left(\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu}+\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu}\right)\ \frac{N(0)}{N(t)}\mathrm{d}t\text{.}$ (6)
-As the objective function depends on the cost, two transformations take place: First, the value $V$ is defined as utility.
-Since the utility function is convex, there is some saturation, resulting in a over time decaying weight $\frac{\partial V(t)}{\partial C(t)}$ applied to the cost.
-Second, the value is discounted.
-For the model of a constant positive discount rate, the discount factor constitutes an exponentially decaying weight $\frac{N(0)}{N(t)}$ applied to the cost.
+The following shows why the unconstrained optimal emission path may result in an unequal distribution of the total cost and the total cost per GDP.
+The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfills (differentiating the objective function (1) with respect to the abatement policy $\mu$)
+[…]
+As the objective function depends on the cost, two transformations take place.
+First, the value $V$ is defined as utility.
+Because the utility function is concave, the weight $\frac{\partial V(t)}{\partial C(t)}$ decays.
+Second, the values are discounted.
+With a constant positive time-preference rate, the discount factor constitutes an exponentially decaying weight $\frac{N(0)}{N(t)}$ applied to the cost.
 The product of both weights represents the cost-to-value-weight.
-The optimal abatement policy $\mu$, representing the equilibrium state of the model, fulfills
-[…]
-From eq.
-7 we see why the model optimization may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered.
-Since $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)>0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)<0$ we see from eq.
-7 that the model equates the increment of the cost of damage to the increment of the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$.
+From eq.˜3, we see that the model optimization may result in some intergenerational inequality: Let $s$ be a fixed time at which a change in the abatement policy $\mu(s)$ is considered.
+Because $\mathrm{d}C_{\mathrm{D}}/\mathrm{d}\mu(s)<0$ and $\mathrm{d}C_{\mathrm{A}}/\mathrm{d}\mu(s)>0$ we see from eq.˜3 that the model equates the decrement in the cost of damage to the increment in the cost of abatement, both weighted by the value-to-cost-sensitivity $\mathrm{d}V(t)/\mathrm{d}C(t)\cdot N(0)/N(t)$.
 We have
-$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{,}$ (8)
-where the to cost function $t\mapsto\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}$ and $t\mapsto\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}$ have different temporal distributions.
-A change in the abatement policy in time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for possibly all $t\geq s$.
-This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e., $\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}$.
-As the model optimization is weighting cost by the cost-to-value weight, we consider the corresponding weighted damage-cost per abatement-cost sensitivity.
-The minus in (9) just accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
-$\rho_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (9)
-From Equation 6 we see that for fixed $s$ the integral over $t\mapsto\rho(s;t)$ equals $1$, i.e., $\rho_{\mathrm{D}}(s)$ can be interpreted as a density (Figure 8).
-Hence, we may interpret $\int_{0}^{\infty}t\cdot\rho_{\mathrm{D}}(s;t)\ \mathrm{d}t$ as the expected time of damage cost per avoided abatement cost in time $s$.
-Figure 8: The density $t\mapsto\rho_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
-In the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement.
-This can be observed in Figure 2, indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity.
-Sensitivity of Cost to Time of Reaching 100 % Abatement
-The sensitivity $\frac{\mathrm{d}}{\mathrm{d}\mu}$ is a functional derivative or gradient.
-To visualize the aggregated dependency structure, we consider the one-parameter deterministic abatement model (5).
-For this abatement model, the equilibrium state is given by
-[…]
-Given the one-parameter model, we may depict the (weighted) damage cost sensitivity $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ and the (weighted) abatement cost sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$, see fig.
-2.
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ -\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\text{,}$ (4)
+where $t\mapsto\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}$ and $t\mapsto\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}$ exhibit different temporal distributions.
+The sensitivity $\frac{\mathrm{d}}{\mathrm{d}\mu}$ is a functional derivative or gradient. Given the one-parameter abatement model (2) the equilibrium state is given by
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ -\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\text{,}$ (5)
+and we can plot the (weighted) damage cost sensitivity $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ and the (weighted) abatement cost sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$, see fig.˜2.
+5.2.4 Damage Cost per Abatement Cost upon Abatement Policy Change
+A change in the abatement policy at time $s$ affects the abatement cost $C_{\mathrm{A}}(s)$ only at time $s$, whereas it affects damage cost $C_{\mathrm{D}}(t)$ for all $t\geq s$.
+This motivates to consider the damage cost per abatement cost induced by a change in the abatement policy $\mu(s)$, i.e.,
+$\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\big{/}\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\text{.}$
+As the model optimization weights cost by the cost-to-value weight, we consider the corresponding weighted damage cost per abatement cost sensitivity.
+Definition (Damage Cost per Abatement Cost upon Abatement Policy Change)
+We define [^2] The minus in (6) accounts for the fact that abatement cost sensitivity and damage cost sensitivity have opposite signs.
+$\gamma_{\mathrm{D}}(s;t)\ :=\ -\frac{\frac{\partial V(t)}{\partial C(t)}\cdot\frac{\partial C_{\mathrm{D}}(t)}{\partial\mu(s)}\cdot\frac{N(0)}{N(t)}}{\frac{\partial V(s)}{\partial C(s)}\cdot\frac{\partial C_{\mathrm{A}}(s)}{\partial\mu(s)}\cdot\frac{N(0)}{N(s)}}\text{.}$ (6)
+Lemma
+At the optimal abatement policy $\mu$, for fixed $s$ the integral over $t\mapsto\gamma_{\mathrm{D}}(s;t)$ equals $1$, i.e., $\gamma_{\mathrm{D}}(s)$ can be interpreted as a density, and we may interpret
+$t_{D}(s)\ :=\ \int_{0}^{\infty}t\cdot\gamma_{\mathrm{D}}(s;t)\ \mathrm{d}t\text{.}$
+as the expected time of damage cost per avoided abatement cost at time $s$.
+The proof is provided in the Appendix B.
+Our numerical analysis will show a significant (expected) time lag for $t_{D}(s)$ (roughly 70 years in a DICE-2016).
+This means that marginal gains of one generation are equated to marginal losses of another generation.
+The optimization does not equalize the burden but rather the weighted marginal burden.
+In this sense, the optimization is indifferent to the temporal distribution of the absolute burden.
 5.3 Model Extensions
-We present modifications to integrated assessment models to investigate the effect of stochastic interest rates, funding of abatement costs, and non-linear financing cost [42].
-These may be also incorporated into more complex IAMs.
-We use the simple DICE model to illustrate potential effects.
+We present modifications to integrated assessment models to investigate the effects of stochastic interest rates, funding abatement costs, and non-linear financing costs [47].
+These may also be incorporated into more complex IAMs.
+We use the DICE model to illustrate potential effects.
 5.3.1 Stochastic Interest Rates
-The classical way in which time-preference rates are modeled in an IAM is via a discount factor.
-A time-$t$ value $V(t)$ is derived in the model, then discounted and aggregated to a final value forming the objective function
-$\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\text{.}$
-We model stochastic interest rates (time-preference rates) $rho$.
-Our implementation allows to use a general discrete forward rate model (LIBOR Market Model).
-Our results were conducted with a classical Hull-White model.
-The model provides the stochastic numeráire $N$ (and thus discount factor) as well as the stochastic forward rate $FR$.
-Setting the volatility of the interest rate to zero recovers the deterministic model with a possibly time-dependent interest rate (term-structure of interest rates).
-For exemplary parametrization, we set the mean reversion speed to $a(t)=0.02$ and the volatility to $\sigma(t)=0.4\%$.
-While our result should be interpreted as a qualitative result, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively.
-For a corresponding visualization, see [40].
-If interest rates are stochastic but the function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM since
-$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (11)
-with $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\right)$.
-Due to the lack of a feedback that allows to adjust the abatement path and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics.
-This changes however, if one assumes that the abatement policy is adapted to the stochastic changes, here the change in interest rate level, i.e., if one allows $\mu$ to be an (adapted) stochastic process.
-[^2] Our choice to make interest rates stochastic through uncertain time preference rates is only exemplary.
-The interest rate is an important factor in linking present abatement cost to the avoided future damage cost.
-While one may introduce stochasticity in many other state variables, the use of an abatement policy adapted to the economic factors (interest rates) will already lead to all other state variables becoming stochastic.
-To model the uncertainty of the rate of time-preference $\rho(t)$, we use a classical Hull-White model [39, 55], that is
-$\begin{split}\mathrm{d}\rho(t)&\ =\ (\theta(t)-a(t)\rho(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad\rho(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ \rho(t)\ \mathrm{d}t\text{,}\end{split}$ (12)
+In classical IAMs, the time preference rate $\rho$ is typically a time-dependent deterministic function, which is often a constant.
+We model a stochastic interest rate (time-preference rate) $\rho$.
+The model provides the stochastic numéraire $N$ (a bank account accruing continuously at rate $\rho$) as well as the stochastic forward rate $FR$.
+Setting the volatility of the interest rate to zero recovers the deterministic model.
+Our implementation allows for the use of a general discrete forward rate model (LIBOR Market Model). Our results were obtained using the classical Hull-White model [48, 49], that is
+$\begin{split}\mathrm{d}\rho(t)&\ =\ (\theta(t)-a(t)\rho(t))\ \mathrm{d}t\ +\ \sigma(t)\ \mathrm{d}W(t)\text{,}\quad\rho(t_{0})\ :=\ r_{0}\text{,}\\ \mathrm{d}N(t)&\ =\ \rho(t)\ N(t)\ \mathrm{d}t\text{,}\end{split}$ (7)
 where an exact time-discretization scheme [50] was used.
 Our stochastic interest rate model is parameterized such that the optimal deterministic abatement strategy would result in the same optimal emission pathway as the model with deterministic interest rates.
+We set $r_{0}$, $\theta(t)$ so that the certainty-equivalent rate $\bar{\rho}(t)$ equals the deterministic time preference rate $\rho$, such that outcomes are exactly identical to the deterministic model (9).
+Mean reversion and volatility were chosen to reflect plausible long-term interest-rate dynamics.
+The chosen mean-reversion speed of $a(t)=0.02$ implies a half-life of $\ln(2)/0.02\approx 35$ years, while the chosen volatility of $\sigma(t)=0.3\%$ produces a dispersion of future rates that remains within a realistic range—roughly $\pm 3$ percentage points around the long-run mean over a century.
+While our result should be interpreted as illustrative rather than predictive, showing that stochastic interest rates generate adverse effects, the interest rate scenarios generated by our parameter choice are realistic, rather conservative: the 1st and 99th percentile of the generated interest rates at t = 100 years are -1 % and 5 %, respectively.
+For a corresponding visualization, see [46].
+As we consider a model extension where $N$ and $V$ are stochastic processes, we redefine the objective function (1) as a random variable. For the optimization we consider the expected welfare
+$\mathrm{E}\left(\int_{0}^{T}V(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t\right)\text{.}$ (8)
+If interest rates are stochastic but the utility function $V(t)$ remains deterministic, adding stochastic interest rates does not change the interaction with the IAM because
+$\begin{split}\mathrm{E}\left(\int_{0}^{T}V(t)\frac{N(0)}{N(t)}\mathrm{d}t\right)&\ =\ \int_{0}^{T}V(t)\ \mathrm{E}\left(\frac{N(0)}{N(t)}\right)\mathrm{d}t\\ &\ =\ \int_{0}^{T}V(t)\exp\left(-\int_{0}^{t}\bar{r}(s)\mathrm{d}s\right)\mathrm{d}t\text{,}\end{split}$ (9)
+where $\bar{r}(t)\ =\ -\frac{\partial}{\partial t}\log\left(\mathrm{E}\left(\exp\left(-\int_{0}^{t}\rho(s)\mathrm{d}s\right)\right)\right)$.
+Owing to the lack of feedback that allows adjustment of the abatement policy and resulting damages, adding stochastic interest rates does not yet introduce changes in the model dynamics.
+The model produces the same emission pathway as for the deterministic model.
+However, this changes if one assumes that the abatement policy is adapted to the (stochastic) changes in the interest rate level.
 5.3.2 Stochastic Abatement Policy
-Assuming that a change in the interest rates triggers a re-calculation of the model optimal abatement policy, a stochastic interest rate implies a stochastic abatement policy, i.e., $\mu$ will become stochastic.
-The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to the changes in interest rates.
+Assuming that a change in the interest rates triggers a re-calculation of the model optimal abatement policy, a stochastic interest rate implies a stochastic abatement policy.
+The abatement policy $t\mapsto\mu(t)$ becomes a stochastic process that adapts to changes in interest rates.
 This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
-The determination of an optimal stochastic abatement policy is a classical optimal exercise problem, as in the valuation of financial products with early exercise options in mathematical finance.
+Determining an optimal stochastic abatement policy is a classical optimal exercise problem, as in the valuation of financial products with early exercise options in mathematical finance.
 Here, $\mu$ takes the role of the exercise strategy.
-As the time-$t$ optimal exercise is required to be a $\mathcal{F}_{t}$-measurable, it may be represented as a functional form of time-$t$ measurable random variables, see [55].
+As the time-$t$ optimal exercise is required to be a $\mathcal{F}_{t}$-measurable, it may be represented as a functional form of time-$t$ measurable random variables [49].
 An approach to determine the optimal exercise strategy is to parameterize $\mu(t)$ as a functional form of $\mathcal{F}_{t}$-measurable random variables and apply a global optimization.
-As we utilize a one-factor Markovian model for the short rate $r$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $\rho(t)$.
-To obtain a robust optimization, reducing the risk of over-fitting or numerical instabilities, we propose a parametric stochastic abatement model. Here, we use the model
+As we utilize a one-factor Markovian model for $\rho$, we consider a functional form where $\mu$ depends on $t$ and the current interest rate level $\rho(t)$.
+To obtain a robust optimization, we propose the parametric stochastic abatement model
 $\mu(t,\omega)=\min\left(\mu_{0}+\left(a_{0}+a_{1}\cdot\rho(t,\omega)\right)\cdot t,1.0\right)\text{,}$ (10)
 where the abatement speed is a (linear) function of the interest rate level $\rho$.
-The model can be seen as a first-order approximation of general policies, where the abatement is allowed to depend in the level of the discount rate $\rho$.
-The simplicity of the model allows us to interpret the parameters.
-We have that $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change of the abatement speed by interest rate change,
+The model can be seen as a first-order approximation of general policies.
+Its simplicity allows the interpretation of the parameters: $a_{0}$ is the abatement speed at the zero interest rate level and $a_{1}$ describes the change in the abatement speed per interest rate change,
 $a_{0}\ =\ \frac{\partial\mu(t)}{\partial t}\big{|}_{r=0}\text{,}\qquad a_{1}\ =\ \frac{\partial}{\partial r}\ \frac{\partial\mu(t)}{\partial t}\text{.}$
-Obviously, the stochastic abatement models may achieve a larger-or-equal welfare or lower-or-equal welfare-risk than the corresponding deterministic model (5) simply due to the additional degree of freedom $a_{1}$ in the abatement optimization.
-The optimization of the model (13) results in a negative value for $a_{1}$, reflecting that higher interest rate correspond to slower abatement speed.
+The optimization of the model (10) results in a negative value for $a_{1}$, reflecting that a higher interest rate corresponds to a slower abatement speed.
 5.3.3 Funding of Abatement Costs
-The geophysical part of the model models a global mean temperature level, which results from past emissions.
 The simple damage function of the DICE model provides damages as a reduction of global GDP.
-Damages may be reduced by performing abatement of emissions.
+Damage may be reduced by abatement of emissions.
 Both abatement and damage are associated with costs.
-The time $t_{i}$ costs are deduced from the time $t_{i}$ GDP.
+The time-$t_{i}$ costs are directly deduced from the time-$t_{i}$ GDP.
 The remainder is then available for consumption or investment.
-Investment adds to the capital which determines the GDP of the next time $t_{i+1}$.
-However, abatement cost and damage cost linked to the same abatement policy decision occur at significantly different times.
-We introduce the option of funding of abatement cost.
-As abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement cost are covered by a loan for which interest corresponds to the current discount rate.
-Our model allows to apply a funding spread, i.
+Investment adds to the capital, which determines the GDP of the next time $t_{i+1}$.
+However, abatement cost and damage cost resulting from the same abatement policy decision occur at significantly different times, as our analysis shows.
+To resolve the time-lag between abatement and damage costs, we introduce the option of funding abatement cost.
+As the abatement of emissions is a planned process of societal relevance, it is reasonable to assume that abatement costs are covered by a loan for which interest corresponds to the current discount rate.
+Our model allows the application of a funding spread, i.
 e.
 higher interest rates for loans than for discounting, but this is not considered in this general introduction of the phenomenon.
-In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments, comparable to the modifications in DICE-2023 [38].
+In future work, the concept of the climate beta could be used to introduce an interest rate spread for abatement investments comparable to the modifications in DICE-2023 [8].
 We define $C_{\mu}(t)$ as the instantaneous abatement costs in time $t$, i.e., the quantity that was formerly denoted by $C_{\mathrm{A}}(t)$.
-We allow that $C_{\mu}(t)$ might be funded for a period $\Delta T_{\mathrm{A}}$.
-Thus abatement costs of $C_{\mu}(t)$ are accrued with the forward rate $FR(t,t+\Delta T_{\mathrm{A}};t)$ observed in $t$, such that the realized abatement cost $C_{\mathrm{A}}$ at time $t+\Delta T_{\mathrm{A}}$ are
+We allow that $C_{\mu}(t)$ is funded for a period $\Delta T_{\mathrm{A}}$.
+Thus, the abatement costs of $C_{\mu}(t)$ are accrued with the forward rate $FR(t,t+\Delta T_{\mathrm{A}};t)$ observed in $t$, such that the realized abatement cost $C_{\mathrm{A}}$ at time $t+\Delta T_{\mathrm{A}}$ is
 $C_{\mathrm{A}}(t+\Delta T_{A})\ :=\ C_{\mu}(t)\ \left(1+FR(t,t+\Delta T_{A};t)\Delta T_{\mathrm{A}}\right)\text{.}$ (11)
 For the case of no funding, i.e. $\Delta T_{\mathrm{A}}=0$, we recover the classical model with $C_{\mu}=C_{\mathrm{A}}$.
-In a standard risk-neutral valuation, this change would have no effect, as the accruing is compensated by the discounting.
-However, the change might introduce an effect in the DICE model, due to the way how abatement and damage cost are associated as well as due to the time-preference included in the utility function.
+Instead of a single (zero-bond like) funding, we can also distribute the time $T_{i}$ cost $C_{\mu}(T_{i})$ to a collection of $m_{i}$ equal value loans with repayments in $T^{k}_{i}$, $k=i+1,\ldots i+m_{i}$.
+In this case the $T_{j}$ cost become
+$C_{\mathrm{A}}(T_{j})\ :=\sum_{i\ :\ \exists\ T^{k}_{i}=T_{j}}\ \frac{1}{m_{i}}\ C_{\mu}(T_{i})\ \left(1+FR(T_{i},T_{j};T_{i})\ (T_{j}-T_{i})\right)\text{.}$ (12)
+From a valuation (discounting) perspective, both funding approaches have no effect, as the accrual is compensated by discounting.
+However, the change in timing might introduce an effect in the DICE model due to the way abatement and damage costs are associated and temporal differences in each generation’s utility or capacity.
+For a fixed emission pathway, the modification leaves present-value abatement costs unchanged (accrual cancels with discounting). The equity effect comes from changes in timing, not value creation.
 Since damages occur instantaneously, we do not consider funding of these. The total cost is given by $C(t):=C_{\mathrm{A}}(t)+C_{\mathrm{D}}(t)$.
+In our numerical experiments, we distributed the funding according to (12) over 60 years ($m_{i}=60$).
 5.3.4 Non-Linear Financing Costs
-For an unsecured financial cash-flow its present value is defined by a discount factor times the cash-flow.
-If the future cash-flow is subject to default, the discount factor is lower, reflecting the additional value reduction due to the risk of (partial) default.
-As default is not an option for future damage cost, it appears as if the risk-free discount factor should apply.
+The present value of an unsecured financial cash-flow is defined by a discount factor times the cash-flow.
+If the future cash-flow is subject to default, the discount factor is lower, reflecting an additional value reduction due to the risk of (partial) default.
+As default is not an option for future damage costs, it appears as if the risk-free discount factor should apply.
 However, since no hedging strategy exists, a risk free funding is not possible.
-Thus additional cost may occur to secure the unsecured funding [42].
+Thus, additional costs may be incurred to secure the unsecured funding [47].
 Since damage cost may become very large - much larger than funds provided by standard financial markets - it is reasonable to assume that these additional funding cost become (over-proportionally) large for larger damages.
 We model this by optionally adding non-linear financing cost using a non-linear funding model [47].
-This model allows, that the discount factor may depend on the magnitude of the cash-flow.
+This model allows the discount factor to depend on the magnitude of the cash-flow.
 Thus the funding of larger cash-flows requires a premium to compensate for a larger default risk or other frictions.
 For our application, this means that larger damages get a larger weight.
-Our model modification is now a modification of the damage cost.
+Consequently, the abatement policy optimization will prefer emissions paths that avoid very large damages, which is beneficial with respect to intergenerational equity.
+We modify the damage cost function.
 Let $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model, i.e.
 the quantity that was formerly denoted by $C_{\mathrm{D}}(t)$.
-We then re-define the effective damage costs as
+We then redefine the effective damage costs as
 $C_{\mathrm{D}}(t)\ =\ C^{\circ}_{\mathrm{D}}(t)\cdot DC(C^{\circ}_{\mathrm{D}}(t);t)\text{,}$
-where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor. It is somewhat similar to the inverse of a discount factor, describing the over-proportional cost to fund large projects.
-As $C^{\circ}_{\mathrm{D}}(t)$ represents a time-value, it is natural that the default compensation factor depends on a normalized value only.
-We allow two different normalizations: either with the numéraire $N(t)$ or with the GDP $Y(t)$.
-Hence, our model for the default compensator is
-$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{N}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{N(t)})\text{,}$ (15)
-or, alternatively,
-$DC(C^{\circ}_{\mathrm{D}}(t);t)\ =\ DC^{Y}(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{Y(t)})\text{.}$ (16)
-The latter approach allows us to penalize damages that exceed a certain percentage of the GDP.
-The factor $DC^{*}(x)$ depends on the size of $x$. For small $x$ we have $DC^{*}(x)=1$, but for large $x$ we may have factors $>1$. Obviously, this will penalize large spikes in the costs.
-In our numerical experiments the funding period was set to 20 years and the non-linear discounting was set to increase the default compensator significantly for damage cost over 3% of the GDP (black).
-The parameter choice is exemplary, other experiments with similar results can be reproduced in the published source code.
+where $DC(C^{\circ}_{\mathrm{D}}(t);t)$ is the default compensation factor modeling over-proportional cost of funding large projects, [47].
+The factor $DC$ is equal to 1 for modest damage values and larger for large damage values.
+As $C^{\circ}_{\mathrm{D}}(t)$ represents a time-value, it is natural that the default compensation factor depends on a normalized value only, e.g., percentage of the GDP.
+In our numerical experiment we used $DC(C^{\circ}_{\mathrm{D}}(t);t)\ :=\ DC^{*}(\frac{C^{\circ}_{\mathrm{D}}(t)}{GDP(t)})$ with $DC^{*}(x)=10\%$ for $x\geq 3\%$ and $DC^{*}(x)=1$ otherwise, basically limiting costs to 3 % of GDP.
 As the abatement cost are usually smaller and part of a more planned process, we do not consider a default compensation factor for the abatement cost.
-4.4.5 Cost per GDP
-To measure the burden born by a generation, we consider the relative total damage and abatement cost per GDP.
-Putting nominal values relative to the GDP is a common measure, e.g., to assess national debt [57].
-Here, it reflects that a more wealthy generation can carry a greater burden, mimicking an effort sharing scheme and also considering the greater capacity of future generations under the expected baseline growth in DICE.
-In our numerical experiments we measure the instantaneous cost-per-GDP $C(s)/GDP(s)$ as well as the lifetime-average cost-per-GDP $\overline{C}(s)/\overline{GDP}(s)$.
-To allow for a netting within a generation one may consider the (discounted) average value over a generation lifetime.
-For a given time $s$ let $T^{\mathrm{L}}(s)$ denote the expected lifetime of a generation.
-We then consider the lifetime average cost and GDP.
-$\overline{C}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}C(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{,}\qquad\overline{GDP}(s)\ :=\ \int_{s}^{s+T^{\mathrm{L}}(s)}GDP(t)\frac{N(s)}{N(t)}\mathrm{d}t\text{.}$
-4.5 Convexity
-Our numerical analysis shows that a continuous adaptation of the abatement strategy may increase the risk and expectation of intergenerational inequality measures as cost-per-GDP.
-The expected emissions/damage/cost are higher than the emissions/damage/cost calculated for the expected interest rate level.
-Such an effect may come as a surprise.
-However, it is well known from mathematical finance, known under the name convexity.
-The effect corresponds to Jensen’s inequality, that the expectation of a convex function of a random variable is higher than the convex function applies to the expectation of the random variable.
-Here, the random variable is the exogenous given discount rate, or stochastic process of the interest rates, and the integrated assessment model acts as a convex function.
-In other words: DICE has convexity in the interest rates once the abatement strategy depends on the interest rate level.
+5.4 Implementation Details
+The numerical experiments were conducted with an annual time-discretization $\Delta t_{i}=1$, using a Monte-Carlo simulation with 10000 sample paths.
+Sensitivities were obtained through adjoint automatic differentiation.
+Optimization was performed with an ADAM optimizer.
 Supplementary Material
 Supplementary material is available online.
-Code and Data availability
-The model code for the stochastic DICE model, including the proposed extensions, as well as the experiments generating the data and figures, is available in the following Git repository https://gitlab.com/finmath/finmath-climate-nonlinear.
 Acknowledgements
@@
 received funding from the German Federal Ministry of Education and Research (BMBF) under the research projects QUIDIC (01LP1907A) as well as by Deutsche Gesellschaft für Internationale Zusammenarbeit (GIZ) GmbH on behalf of the Government of the Federal Republic of Germany and Federal Ministry for Economic Cooperation and Development (BMZ).
-Author Contribution
+Author Contributions
 C.F.
@@
 The authors declare that they have no competing interests.
+[…]
+[…]
 Supplementary information for
 Intergenerational Equitable Climate Change Mitigation:
-Negative Effects of Stochastic Interest Rates;
+Negative Effects of Stochastic Interest Rates and;
 Positive Effects of Financing
+Christian P. Fries and Lennart Quante
+Contents
 […]
+[supplement]l1
+Appendix A Supplementary Figures
+We provide some additional figures.
+LABEL:fig:cost_abatement_damage_disounted shows the temporal distribution of the cost, which should be compared to the temporal distribution of the cost sensitivity in fig.˜2.
+LABEL:fig:pathways_deterministic_with_extension and LABEL:fig:pathways_stochastic_with_extension show the associated model quantities (emission, abatement, damage, and total cost per GDP) for the deterministic and stochastic model, with and without our model extensions.
+LABEL:fig:opt_abatement_time_by_interest_level and LABEL:fig:opt_abatement_time_by_interest_volatility show the sensitivity of the abatement policy on the interest rate level and interest rate volatility.
 […]
-[…]
-[…]
-figureSupplementary figures
-Variation of the Interest Rate Volatility
-Supplementary Figure 1: Stochastic DICE Model with Different Interest Rate Volatilities ($\sigma=0.2\%$ (red), $\sigma=0.3\%$ (blue), $\sigma=0.4\%$ (green)).
+figureList of Supplementary Figures
+Supplementary Figure 1: The abatement cost $C_{\mathrm{A}}$ (green) and damage cost $C_{\mathrm{D}}$ (red), discounted by $\frac{N(0)}{N(t)}$, in the original DICE model.
+The optimization is not balancing the temporal distribution of the cost - which is ultimately resulting in an intergenerational inequality.
+supfigure]fig:cost_abatement_damage_disounted
+Supplementary Figure 2: Deterministic model: Financing extensions lead to faster abatement compared to standard model; a emissions, b abatement, c damages, and d total GDP relative cost.
+supfigure]fig:pathways_deterministic_with_extension
+Supplementary Figure 3: Stochastic model: Financing extensions lead to faster abatement compared to standard model; a emissions, b abatement, c damages, and d total GDP relative cost.
+Shaded area shows the 10th to 90th percentiles (VaR).
+supfigure]fig:pathways_stochastic_with_extension
+Supplementary Figure 4: Stochastic DICE Model with Different Interest Rate Volatilities ($\sigma=0.2\%$ (green), $\sigma=0.3\%$ (blue), $\sigma=0.4\%$ (red)).
 Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost).
 Shaded area shows the 10th to 90th percentiles.
-Variation of the Interest Rate Level
+supfigure]fig:pathways_stochastic_variation
 Supplementary Figure 5: Increasing interest rate level increases the time of 100% abatement $T^{\mu=1}$.
+supfigure]fig:opt_abatement_time_by_interest_level
 Supplementary Figure 6: Increasing interest rate volatility, i.e. riskiness, decreases the time of 100% abatement $T^{\mu=1}$.
+supfigure]fig:opt_abatement_time_by_interest_volatility
+Appendix B Damage Cost per Abatement Cost upon Abatement Policy Change
+We show that the damage cost per abatement cost upon abatement policy change is a density if $\mu$ is in its equilibrium state.
+If $\mu$ denotes the optimal abatement policy we have
+[…]
+For the time discrete model where $C_{\mathrm{A}}(s)$ depends on $\mu(s)$ only, we find at $s=t_{i}$ that
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ =\ \frac{\mathrm{d}V(t_{i})}{\mathrm{d}C(t_{i})}\frac{\mathrm{d}C_{\mathrm{A}}(t_{i})}{\mathrm{d}\mu(t_{i})}\ \frac{N(0)}{N(t_{i})}\ \Delta t_{i}\text{.}$
+Hence, for $\Delta t_{i}=1$ (using $s=t_{i}$)
+$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ =\ -\frac{\mathrm{d}V(s)}{\mathrm{d}C(s)}\frac{\mathrm{d}C_{\mathrm{A}}(s)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(s)}\text{.}$
+and thus
+$\int_{0}^{T}-\frac{\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}}{\frac{\mathrm{d}V(s)}{\mathrm{d}C(s)}\frac{\mathrm{d}C_{\mathrm{A}}(s)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(s)}}\ \mathrm{d}t\ =\ 1\text{.}$
+Appendix C Non-Linear Financing Costs
+We modify the damage cost function as
+$C_{\mathrm{D}}(t)\ =\ C^{\circ}_{\mathrm{D}}(t)\cdot DC(C^{\circ}_{\mathrm{D}}(t);t)\text{,}$
+where $C^{\circ}_{\mathrm{D}}(t)$ denote the damage cost of the classical model and $DC(C^{\circ}_{\mathrm{D}}(t);t)$ the default compensation factor modeling the over-proportional cost of funding large projects, [47].
+We allow for two different normalizations: with the numéraire $N(t)$ or with the GDP $Y(t)$. Given a function $DC^{*}$, the model for the default compensator is
+$DC\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ =\ DC^{N}\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ :=\ DC^{*}\left(\frac{C^{\circ}_{\mathrm{D}}(t)}{N(t)}\right)\text{,}$ (13)
+or, alternatively,
+$DC\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ =\ DC^{Y}\left(C^{\circ}_{\mathrm{D}}(t);t\right)\ :=\ DC^{*}\left(\frac{C^{\circ}_{\mathrm{D}}(t)}{Y(t)}\right)\text{.}$ (14)
+The latter approach allows us to penalize damages that exceed a certain percentage of the GDP.
+The factor $DC^{*}(x)$ depends on the size of $x$. For small $x$ we have $DC^{*}(x)=1$, but for large $x$ we may have factors >1>1. Obviously, this will penalize large spikes in the costs.
+Appendix D Model Implementation
+We provide a complete open-source implementation [51] of the model illustrating all our extensions, utilizing the Monte-Carlo implementation of standard interest rate models for simulation and (stochastic) algorithmic differentiation to efficiently calculate the sensitivities of model quantities against each other.
+Our implementation can deal with an arbitrary time-discretization $\{t_{i}\}$ and the DICE model is redefined as the Euler-discretization of a time-continuous model.
+For the equidistant discretisation with $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ the classical DICE-2016 model is recovered.
+Our numerical experiments are based on a time-discretization with $\Delta t_{i}=1\leavevmode\nobreak\ \text{year}$, however, the differences to the $\Delta t_{i}=5\leavevmode\nobreak\ \text{years}$ discretization of the classical DICE-2016 are negligible in our study.
+With annual time-discretization the piecewise constant abatement policy constitutes $499$ free parameters for a $T=500\leavevmode\nobreak\ \text{years}$ model.
+In the numerical experiments an annual time-discretization $\Delta t_{i}=1$ was
+Appendix E Associated Concepts
+E.1 Time Preference Rates and the Ramsey Equation
+The DICE model extends a standard neoclassical optimal growth model (the Ramsey model) with economic impacts from climate change.
+We give a short recapitulation of the Ramsey equation since it makes clear that our modelling of a stochastic time preference rate $\rho$ is equivalent to a stochastic social discount rate $r$.
+Given that the utility $U$ is a function of the time-$t$ consumption $C(t)$, the discounted marginal utility is
+$U^{\prime}(C(t))\cdot\exp(-\rho t)\text{.}$
+If we assume a declining marginal utility (constant relative risk aversion)
+$\mathrm{d}U^{\prime}/U^{\prime}\ =\ -\eta\ \mathrm{d}C/C\text{,}$
+where $\eta$ denotes the elasticity of marginal utility, then we have
+$\frac{\mathrm{d}}{\mathrm{d}t}\left(U^{\prime}(C(t))\cdot\exp(-\rho t)\right)$
+$\ =\ -\eta\cdot U^{\prime}(C(t))\cdot\frac{\mathrm{d}C(t)}{\mathrm{d}t}/C(t)\cdot\exp(-\rho t)-\rho\cdot U^{\prime}(C(t))\cdot\exp(-\rho t)$
+$\ =\ -(\eta\cdot g+\rho)\cdot U^{\prime}(C(t))\cdot\exp(-\rho t)\ =\ -r\cdot U^{\prime}(C(t))\cdot\exp(-\rho t)\text{.}$
+with
+$r\ =\ \eta\cdot g+\rho\text{,}$ (15)
+where $g=\frac{\mathrm{d}C(t)}{\mathrm{d}t}/C(t)=\frac{\mathrm{d}}{\mathrm{d}t}\ln(C(t))$. In this context, $r$ is called the social discount rate and Equation (15) is the Ramsey equation.
+The (optimal) growth of the consumption $g$ depends on the optimal emission and savings paths, and is an outcome of the model optimization.
+The elasticity of marginal utility $\eta$ is an input to the model and usually empirically estimated [52].
+For the link between time preference and social discount rate see also [13].
+The time-preference $\rho$ is an exogenous parameter and subject to controversial discussions.
+That $\rho$ is perceived as exogenous is already transparent from Nordhaus 2017 paper on the DICE model [7], showing a dependency of the social cost of carbon on the chosen (growth-adjusted) discount rate.
+In [53] an analysis is performed on the dependency of the social cost of carbon on a distribution of the discount rate. Here, the Bond market interest rate is used as an input to the model.
+Via the (15) we see with a constant given that $\eta$ a stochastic time preference rate $\rho$ corresponds to a stochastic social discount rate $r$.
+While [53] considers a distribution of discount rates derived from Bond market rates, there is no feedback effect between the integrated assessment model and the given interest rates.
+The novelty of our extensions is to consider a stochastic process of future uncertain discount rates and consider its interaction with an integrated assessment model.
+This models the possibility that policymakers adapt the abatement strategy to changes in interest rates.
+Such an adaptation is plausible, because already the static analysis shows that the optimal abatement strategy depends on the level of the interest rate.
+E.2 Social Cost of Carbon
+A prominent example of a metric describing the temporal structure of cost is the social cost of carbon (SCC).
+Model extensions updating IAMs [12, 13, 20, 54] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [55] or energy consumption [56], but also general SCC [43] emphasize the need to improve the precision of the estimated SCC, especially by including additional impact channels.
+The social cost of carbon is defined as
+$SCC(t)\ :=\ -1000\cdot\frac{\partial V(t)}{\partial E(t)}\Big{/}\frac{\partial V(t)}{\partial C(t)}\text{.}$ (16)
+The $SCC$ is the time-$t$ marginal cost of emitting one additional unit of carbon.
+Its unit is $SCC]=\frac{\mathrm{USD}}{\mathrm{tCO}_{2}}$.
+We would like to note that depicting $SCC$ in the form (16) may be misleading, since it is an undiscounted time-$t$ value.
+Instead, one should rather consider the discounted, or, numéraire relative $SCC^{N}$, i.e.,
+$SCC^{N}(t)\ :=\ -\frac{1000}{N(t)}\cdot\frac{\partial V(t)}{\partial E(t)}\Big{/}\frac{\partial V(t)}{\partial C(t)}\text{.}$
+The “social cost of carbon” is maybe not a suitable measure to access the temporal distribution of the social burden, and hence intergenerational equity, because pricing emissions at the level of the SCC does not cover the cost associated with the climate mitigation paths: the accumulated value $\int_{0}^{T}SCC(t)\cdot E(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$ does not match the accumulated cost $\int_{0}^{T}C(t)\cdot\frac{N(0)}{N(t)}\mathrm{d}t$, see [45].